Jefferson Lab HALL A Logbook for May 09 - All EXCEPT Automatic Entries

  • Main INDEX

    Sunday

  • 278017 05/31/09 22:25 dwang NMR 63.07%
  • 278003 05/31/09 20:44 Boeglin New HAPPEX run 31453
  • 278000 05/31/09 20:21 adaq Hall A Tools
  • 277999 05/31/09 20:20 Boeglin Spot ++
  • 277973 05/31/09 18:34 adaq ROC3 Graphs
  • 277971 05/31/09 18:22 dwang NMR 62.198%
  • 277951 05/31/09 16:01 ellie RHRS Q1 Red Helium Alerts during Day Shift
  • 277950 05/31/09 16:01 ellie RHRS Q1 Red Helium Alerts during Day Shift
  • 277949 05/31/09 16:01 ellie Day Shift Summary
  • 277948 05/31/09 15:56 ellie Oops, Run # was 31451
  • 277943 05/31/09 15:42 flay BB MWDC (Run# 3011)
  • 277942 05/31/09 15:32 flay Neutron Detector TDCs (Run# 21889)
  • 277939 05/31/09 15:26 ellie Started New HAPPEX Run # 31452
  • 277932 05/31/09 14:40 flay RHRS TDCs Run# 21887
  • 277927 05/31/09 13:58 luhj NMR62.4%
  • 277922 05/31/09 13:36 R. Michaels remove adaql3 halogscript cron / doc of failure mech.
  • 277917 05/31/09 13:16 ellie VDC Efficiencies Screen Grabs
  • 277916 05/31/09 13:10 ellie RHRS Dipole Helium Level Low
  • 277911 05/31/09 12:38 ellie Hall A Tools Screen Grab
  • 277910 05/31/09 12:37 ellie Spot++ Screen Grab
  • 277909 05/31/09 12:34 flay Run# 3005 Deadtime
  • 277904 05/31/09 12:18 flay EPICS Variables Run# 3004
  • 277886 05/31/09 11:07 ellie Started New HAPPEX Run # 15263
  • 277881 05/31/09 10:52 ellie HAND N3-3R Looks good again
  • 277880 05/31/09 10:51 flay EPICS Variables
  • 277879 05/31/09 10:49 flay High Deadtime for Run# 3001
  • 277878 05/31/09 10:36 ellie HAND N3-3R ADC & TDC low for Run 21878
  • 277870 05/31/09 09:54 luhj run3000_3H3_pol61.6
  • 277869 05/31/09 09:53 luhj run3000_3H3_pol61.6
  • 277868 05/31/09 09:48 ellie Started New Compton Run # 17099
  • 277863 05/31/09 09:14 ellie Avian Visitor
  • 277862 05/31/09 09:00 ellie Low Helium Flow Red Alarm on Right Q1 again
  • 277857 05/31/09 08:31 ellie Both LHRS & RHRS DAQs Crashed, rebooted, taking data
  • 277856 05/31/09 08:29 ellie RHRS Q1 Giving Red Helium Flow Alarm
  • 277846 05/31/09 08:15 ellie RHRS Q1 Heling flow low, giving yellow warning
  • 277844 05/31/09 08:00 miham OWL Shift Summary (05/31/09)
  • 277839 05/31/09 07:43 miham New Compton Run #17098
  • 277823 05/31/09 05:57 posik target strip charts
  • 277822 05/31/09 05:56 posik Laser absorption spectrum
  • 277821 05/31/09 05:53 miham Compton now working - New Run #17097
  • 277820 05/31/09 05:40 posik NMR 61.6%
  • 277815 05/31/09 05:19 miham Q1 alarms
  • 277814 05/31/09 05:15 miham Spot++ check (#2990)
  • 277788 05/31/09 03:38 miham Funny Compton - New Compton run #17098
  • 277773 05/31/09 01:50 miham Compton problem
  • 277770 05/31/09 01:21 posik NMR 61.3%
  • 277761 05/31/09 00:34 miham Spot++ check (#2980)
  • 277760 05/31/09 00:30 miham BB, ND, HRS-R HV GUI screen shot - All green
  • 277754 05/31/09 00:03 Boeglin Shift Summary
  • Saturday

  • 277733 05/30/09 21:54 adaq Hall A Tools Screen
  • 277730 05/30/09 21:30 Jin Huang NMR 56.8%
  • 277726 05/30/09 21:25 Jin Huang Change Calibration Constant; EPR Calibration for Longitudinal- Field of Cell Moss; Correct Coils has been ON
  • 277724 05/30/09 21:25 Boeglin Spot ++ after beam back
  • 277715 05/30/09 20:24 Jin Huang NMR 58.7%
  • 277705 05/30/09 20:09 R. Michaels dont run 2 halogscripts
  • 277704 05/30/09 19:58 R. Michaels halogscript
  • 277703 05/30/09 19:46 A. Camsonne Put halogscript in crontab on adaql3
  • 277702 05/30/09 19:39 adaq half wave plate change
  • 277685 05/30/09 17:19 A. Deur NMR polarimetry: 58.5%
  • 277676 05/30/09 16:17 Boeglin Correction for Q1 probelm: it is the Right Q1
  • 277673 05/30/09 16:13 adaq Left Q1 LN level low
  • 277672 05/30/09 16:08 A. Deur absorption spectrum
  • 277670 05/30/09 16:00 S. Sirca Day shift SUMMARY
  • 277665 05/30/09 15:35 S. Sirca Raw rates snapshot
  • 277652 05/30/09 14:22 S. Sirca Hall A General Tools
  • 277651 05/30/09 14:22 S. Sirca Spot++ check with run 2960
  • 277642 05/30/09 13:01 Norum NMR: Pol = 59.5%
  • 277640 05/30/09 13:00 S. Sirca Started new HAPPEX run #31446
  • 277619 05/30/09 11:55 S. SIrca Compton HV tripped && restored
  • 277595 05/30/09 09:08 Holmstrom ND N3 right
  • 277590 05/30/09 08:42 ellie Shift Checklist Website
  • 277587 05/30/09 08:14 S. Sirca Right HRS Q1 entry in ELOG
  • 277586 05/30/09 08:04 posik NMR 58.5%
  • 277584 05/30/09 08:04 Ibrahim Owl Shift Summary (May 30, 2009)
  • 277581 05/30/09 08:01 Ibrahim Low Right Q1 Flow Alarm Every Few Minutes!
  • 277580 05/30/09 07:57 eip BB MWDC
  • 277575 05/30/09 07:06 Ibrahim Restarted the HAPPEX DAQ. New Run 31445
  • 277570 05/30/09 06:49 Ibrahim Low Right Q1 Flow
  • 277569 05/30/09 06:44 Ibrahim Restarted the Compton DAQ. New Run 17085
  • 277560 05/30/09 05:59 Ibrahim Checklist Sheets
  • 277559 05/30/09 05:47 posik target strip charts
  • 277558 05/30/09 05:46 posik Laser absorption spectrum
  • 277553 05/30/09 05:36 miham New golden root files #2942 and #21819
  • 277552 05/30/09 05:27 Ibrahim Spot++ for Run 2942
  • 277547 05/30/09 05:07 Ibrahim Rates for He3 cell at 4x4 mm, 2 uA and BPMB x = -0.5 mm
  • 277542 05/30/09 05:00 Ibrahim Rates for He3 cell at 6x6 mm, 2 uA and BPMB x = -0.5 mm
  • 277533 05/30/09 04:49 Ibrahim Rates He3 cell 6x6 mm 2 uA BPMB x = -1.0 mm
  • 277522 05/30/09 04:14 posik NMR 57.1%
  • 277508 05/30/09 03:53 posik target->optics
  • 277505 05/30/09 03:22 Ibrahim Restarted the HAPPEX DAQ. New Run 31444
  • 277500 05/30/09 02:48 Ibrahim Recent Beam Position Chart
  • 277499 05/30/09 02:36 Ibrahim Spot++ for Run 2933
  • 277494 05/30/09 02:25 Ibrahim Rates 4x4 mm, 2 uA, BPMB x = -1.0
  • 277485 05/30/09 02:10 A. Camsonne Rates Deuterium 2 uA 4x4 mm BPMB X at -1.5 mm
  • 277482 05/30/09 02:06 Jin Huang Check Laser Fiber Temperature
  • 277474 05/30/09 02:01 A. Camsonne Rates deuterium 2 uA 4x4 mm BPMB X to 0.5
  • 277472 05/30/09 01:58 A. Camsonne Reference cell centering
  • 277468 05/30/09 01:54 A. Camsonne Rates deuterium 2uA 4x4 mm
  • 277464 05/30/09 01:49 Jin Huang Reducing Transverse Pickup Coil Pre-amplification
  • 277463 05/30/09 01:46 A. Camsonne Rates deuteriums 2uA Raster 6x6 mm
  • 277456 05/30/09 01:37 A. Camsonne Deuterium Ref cell 5 uA rates
  • 277450 05/30/09 01:24 A. Camsonne Target motion trouble fixed
  • 277447 05/30/09 00:39 Jin Huang Target Ladder Stuck
  • 277446 05/30/09 00:33 Jin Huang Spin Rotated to Long+ (0 degree)
  • 277444 05/30/09 00:09 Boeglin Shift Summary
  • Friday

  • 277441 05/29/09 22:58 A. Camsonne Scalers He3 11uA
  • 277432 05/29/09 21:58 adaq HallA Tools Screen
  • 277431 05/29/09 21:56 Yawei Fill D2 into reference cell, 135psig
  • 277430 05/29/09 21:56 Boeglin Spot++
  • 277429 05/29/09 21:49 Norum NMR: Pol = 59.1%
  • 277424 05/29/09 21:16 A. Camsonne Compton new decoder in production
  • 277406 05/29/09 20:00 hoyoung Right arm VDC efficiency
  • 277397 05/29/09 18:53 ellie HAND HV Change
  • 277388 05/29/09 17:14 Yawei Neutron detector N1 plane left 0 ADC and TDC
  • 277383 05/29/09 16:39 Jin Huang Continue: Target: Quarter Wave Plate Optimization Tests
  • 277382 05/29/09 16:35 Norum NMR: Pol = 58.0%
  • 277376 05/29/09 16:00 X. Zheng day shift summary
  • 277368 05/29/09 14:45 A. Camsonne Compton electron detector in
  • 277363 05/29/09 14:21 X. Zheng started HAPPEX run 31442
  • 277362 05/29/09 14:04 X. Zheng Beam HWP inserted at 11:52
  • 277354 05/29/09 13:58 X. Zheng spot++ check
  • 277352 05/29/09 13:27 R. Michaels NMR 58.1
  • 277351 05/29/09 13:27 X. Zheng Beam is back
  • 277350 05/29/09 13:10 X. Zheng stopped HAPPEX run 31441
  • 277349 05/29/09 12:48 X. Zheng injector optimization (estimate 1 hour)
  • 277346 05/29/09 12:44 R. Michaels Tgt spin state -- ref histos
  • 277332 05/29/09 12:03 X. Zheng started HAPPEX run 31441
  • 277329 05/29/09 11:54 R. Michaels He3 tgt put in at 10:21 (LATE entry)
  • 277328 05/29/09 11:52 X. Zheng stopped HAPPEX run 31440
  • 277321 05/29/09 11:22 ellie HAND Golden Run set to 21779 to reflect ADC fix
  • 277320 05/29/09 11:14 R. Michaels magnets seem stable now
  • 277315 05/29/09 10:51 R. Michaels L-HRS DAQ reboot
  • 277313 05/29/09 10:44 X. Zheng this run has fluctuating left dipole
  • 277308 05/29/09 10:41 fansler Left Dipole cryo fault
  • 277307 05/29/09 10:38 fansler Q1 cryo problem
  • 277306 05/29/09 10:34 X. Zheng Left dipole update
  • 277295 05/29/09 10:22 R. Michaels L-HRS dipole He fault
  • 277294 05/29/09 10:21 R. Michaels NMR 56.6
  • 277291 05/29/09 10:17 X. Zheng spot++ on BeO
  • 277288 05/29/09 10:07 X. Zheng Back to beam permit
  • 277283 05/29/09 09:08 Higinbotham N2-5L Fixed (signal left in scope)
  • 277282 05/29/09 08:44 R. Michaels target to BeO
  • 277281 05/29/09 08:38 X. Zheng stopped HAPPEX run 31439
  • 277280 05/29/09 08:34 A. Camsonne Asked for controlled access to check magnets
  • 277279 05/29/09 08:07 X. Zheng Q1 (right) tripped again, requesting for controlled access
  • 277277 05/29/09 07:57 ejensen OWL shift summary
  • 277276 05/29/09 07:43 A. Camsonne TECHS : Q1 tripped and get cyrogenic fault when ramping up
  • 277275 05/29/09 07:20 A. Camsonne Q1s tripped again
  • 277274 05/29/09 07:18 ejensen Q1 trips (again)
  • 277271 05/29/09 07:06 posik Laser absorption spectrum
  • 277268 05/29/09 06:28 posik NMR Scan 58.2%
  • 277263 05/29/09 05:37 A. Camsonne Right Q1 back
  • 277261 05/29/09 05:23 A. Camsonne Right Q1 cryo fault again
  • 277260 05/29/09 05:20 A. Camsonne No problem with cryo, ramping up left and right Q1
  • 277258 05/29/09 05:00 A. Camsonne Cryo levels low
  • 277246 05/29/09 03:57 posik Laser absorption spectrum
  • 277245 05/29/09 03:52 posik target strip charts
  • 277232 05/29/09 01:55 posik NMR 58.3%
  • 277223 05/29/09 01:03 ejensen MF2 alarm
  • 277222 05/29/09 00:38 eip BB MWDC
  • Thursday

  • 277216 05/28/09 23:55 Jin Huang Shift Summary - May 28th Swing
  • 277207 05/28/09 23:00 M.Shabestari Target strip charts
  • 277202 05/28/09 22:24 M.Shabestari Laser absorption spectrum
  • 277197 05/28/09 21:49 A. Camsonne Compton photon asymmetry assymetrical
  • 277196 05/28/09 21:38 Jin Huang Continue: Target: Quarter Wave Plate Optimization Tests
  • 277193 05/28/09 21:19 M.Shabestari NMR frequency sweep
  • 277190 05/28/09 21:12 Jin Correction: Spin Rotated to Transverse- (270 degree) during down time
  • 277189 05/28/09 21:03 Jin Huang Hall A Tools, Swing 28th
  • 277188 05/28/09 20:56 M.Shabestari Target strip charts
  • 277183 05/28/09 20:42 A. Camsonne Compton vertical scan
  • 277182 05/28/09 20:38 A. Camsonne Beam position during vertical scan
  • 277181 05/28/09 20:11 M.Shabestari Laser absorption spectrum
  • 277167 05/28/09 18:30 ellie HAND HV: It's there!, ND Golden Run Set to 21757
  • 277166 05/28/09 18:16 ellie HAND: HV gettin' there
  • 277161 05/28/09 17:50 Yawei Happex plots after changing the HV
  • 277160 05/28/09 17:35 ellie HAND HV: Trying to turn the Bad to Good
  • 277159 05/28/09 17:33 ellie HAND: The Good, the Bad, and the Ugly
  • 277156 05/28/09 17:31 Jin Huang Happex Run Restarted
  • 277153 05/28/09 17:27 M.Shabestari Target Strip charts
  • 277152 05/28/09 17:17 hoyoung L-arm Cerenkov ADCs 6
  • 277151 05/28/09 17:16 Yawei Lumi HV changed
  • 277150 05/28/09 17:14 M.Shabestari Laser absorption spectrum
  • 277147 05/28/09 16:50 Jin Huang Trigger rate+Spot Check @ He3, 6x6 raster
  • 277146 05/28/09 16:47 M.Shabestari NMR frequency sweep
  • 277142 05/28/09 16:41 Jin Huang Spot check at He3; 4x4 raster
  • 277140 05/28/09 16:40 M.Shabestari Target--->3He
  • 277134 05/28/09 16:32 Jin Huang Spot++ @ Carbon, 7x5mcc
  • 277131 05/28/09 16:28 Jin Huang BeO Camshot
  • 277130 05/28/09 16:27 Jin Huang Harp Scan Result
  • 277129 05/28/09 16:26 Moffit ROC5 checks today...
  • 277124 05/28/09 16:03 S. Sirca Day shift SUMMARY
  • 277122 05/28/09 15:45 X. Zheng more details
  • 277121 05/28/09 15:42 A. Camsonne HV GUI restarted
  • 277120 05/28/09 15:41 A. Camsonne Compton electron detector check
  • 277119 05/28/09 15:35 ellie HAND Access Activities, HV Changed
  • 277118 05/28/09 15:32 adev BigBite WireChamber Measured Voltage change!
  • 277117 05/28/09 15:26 Ramesh/Xiaoc Access activity [PVDIS work]
  • 277116 05/28/09 15:22 adev BigBite WireChamber Measured Voltage change!
  • 277115 05/28/09 15:14 A. Camsonne Compton electron detector out
  • 277112 05/28/09 14:51 R. Michaels re: hatsv3 problems
  • 277109 05/28/09 14:09 miham New width of the T5 trigger pulse
  • 277098 05/28/09 12:50 R. Subedi hatsv3 (RHRS HV portserver) power cycled
  • 277093 05/28/09 12:35 averett Reference cell repacel H2 with D2
  • 277088 05/28/09 12:23 A. Camsonne Hall A in controlled access Full survey done
  • 277074 05/28/09 12:03 Jin Huang Spin Rotated to Transverse- (90 degree) during down time
  • 277058 05/28/09 11:14 S. Sirca Beam current excursion
  • 277055 05/28/09 10:53 S. Sirca Beam will be off at 11:20
  • 277054 05/28/09 10:53 S. Sirca General Tools shortly before beam off
  • 277053 05/28/09 10:47 S. Sirca Spot++ check just before beam off
  • 277048 05/28/09 09:59 Techs/spiege Gas bottles
  • 277039 05/28/09 08:05 miham OWL Shift Summary (05/28/09)
  • 277037 05/28/09 07:48 miham NMR frequency sweep (polarization 58.5% ) and Laser absorption spectrum.
  • 277032 05/28/09 07:40 miham Another Compton HV trip
  • 277031 05/28/09 07:33 miham New Compton run #17050
  • 277030 05/28/09 07:13 adaq Left Arm Cerenkov ADCs
  • 277025 05/28/09 06:42 R. Subedi Re: HV gui for BB, HRS-R, ND, LUMI
  • 277020 05/28/09 06:11 miham Target Strip Charts
  • 277015 05/28/09 05:34 miham New Compton run #17049
  • 277010 05/28/09 04:58 miham Target strip charts
  • 277009 05/28/09 04:57 miham Spot++ check (#2839)
  • 277004 05/28/09 04:17 miham Reference cell pressure gauge calibration
  • 276990 05/28/09 03:20 miham New Compton run #17048
  • 276985 05/28/09 02:59 Jin Huang Target: Quarter Wave Plate Optimization Tests
  • 276983 05/28/09 02:39 miham NMR Frequency Sweep , polarization 57.8%
  • 276979 05/28/09 02:30 miham Compton Strip charts after the HV trip
  • 276978 05/28/09 02:27 miham Hall-A ALARM - Compton HV trip
  • 276973 05/28/09 01:46 miham ROC 5 PROBLEMS - follow up
  • 276972 05/28/09 01:31 miham ROC 5 PROBLEMS
  • 276968 05/28/09 01:01 miham Scaler snap shots (#2834)
  • 276967 05/28/09 00:56 miham New Compton Run #17046 - Had some problems before new run
  • 276966 05/28/09 00:45 miham Spot++ check (#2834)
  • 276965 05/28/09 00:41 miham Laser absorption spectrum 05-27-09 (00:43)
  • 276964 05/28/09 00:39 miham HV gui for BB, HRS-R, ND, LUMI
  • 276963 05/28/09 00:35 miham Target Strip charts
  • 276957 05/28/09 00:00 X. Zheng swing shift summary
  • Wednesday

  • 276952 05/27/09 23:04 X. Zheng Should be HAPPEX run 31432
  • 276951 05/27/09 23:01 X. Zheng Start HAPPEX run 31433
  • 276946 05/27/09 22:49 Yawei Roc3 sync gaps,run#2831
  • 276945 05/27/09 22:26 pkpan NMR_58%
  • 276940 05/27/09 21:58 A. Camsonne Continue running in those conditions
  • 276939 05/27/09 21:58 A. Camsonne Continue running in those conditions
  • 276934 05/27/09 21:25 John Watson N1-N5 TDC Problems gone
  • 276929 05/27/09 20:50 X. Zheng spot++ run 2828
  • 276928 05/27/09 20:29 yawei Lumi is saturated
  • 276910 05/27/09 19:22 X. Zheng Started HAPPEX run 31431 at 18:40
  • 276901 05/27/09 18:29 adaq TDC spectra for N1-N5
  • 276900 05/27/09 18:12 R. Michaels Plot of Moller polarizations.
  • 276897 05/27/09 18:04 pkpan NMR_59%
  • 276892 05/27/09 17:58 X. Zheng 6x6 raster on pol 3He
  • 276889 05/27/09 17:55 A. Camsonne, BB HV back on
  • 276880 05/27/09 17:41 X. Zheng spot++ on BeO
  • 276877 05/27/09 17:20 adev BigBite WireChamber Measured Voltage change!
  • 276876 05/27/09 17:19 adev BigBite WireChamber Measured Voltage change!
  • 276875 05/27/09 17:05 R. Michaels IHWP in/out effect
  • 276874 05/27/09 16:57 R. Michaels re: Moller result
  • 276872 05/27/09 16:50 A. Saha Moller Measurements
  • 276872 05/27/09 16:42 A. Saha Moller Measurements
  • 276871 05/27/09 16:40 A. Camsonne portserver hatsv3 stuck
  • 276870 05/27/09 16:11 tireman Day Shift Summary
  • 276868 05/27/09 15:25 A. Camsonne Asked for Half Wave Plate out
  • 276867 05/27/09 13:48 adev BigBite WireChamber Measured Voltage change!
  • 276866 05/27/09 13:42 A. Camsonne HVGUI killed on adaql2
  • 276865 05/27/09 13:37 adev BigBite WireChamber Threshold change!
  • 276864 05/27/09 13:26 Moffit hvgui running on adaql2.. should switch to a different machine
  • 276863 05/27/09 13:12 tireman fire alarm test
  • 276862 05/27/09 12:56 A. Camsonne Moeller measurement procedure starting
  • 276861 05/27/09 12:52 averett target to Empty Pos'n
  • 276860 05/27/09 12:51 tireman HAPPEX run end for moeller measurement
  • 276855 05/27/09 12:10 D. Parno At next Compton restart, please switch to Prod_FT configuration
  • 276843 05/27/09 11:59 tireman scaler snap shots at 11 uA beam
  • 276842 05/27/09 11:54 tireman DAQ struggles at steady beam
  • 276841 05/27/09 11:44 tireman Hall A tools and BCM
  • 276840 05/27/09 11:42 tireman spot++ check for shift checklist
  • 276839 05/27/09 11:35 tireman Started New Happex run # 31430
  • 276838 05/27/09 11:29 tireman Beam current change to 11 uA
  • 276835 05/27/09 11:23 averett move to 3He target, NMR measurement
  • 276832 05/27/09 11:14 tireman Run 2815 no run start entry
  • 276831 05/27/09 10:56 averett N2 in reference cell
  • 276830 05/27/09 10:56 tireman scaler snap shots during N2 reference cell run
  • 276829 05/27/09 10:44 tireman Runs 2815 and 21715 are suspect end early for Left HRS DAQ
  • 276820 05/27/09 10:05 tireman H2 Reference cell scaler rates
  • 276819 05/27/09 10:02 tireman New HAPPEX run started
  • 276816 05/27/09 09:34 averett move to ref cell
  • 276809 05/27/09 08:11 averett NMR Measurement
  • 276803 05/27/09 08:02 miham OWL Shift Summary (05/27/09)
  • 276802 05/27/09 07:59 miham New Golden runs #2810 and #21710 after the Current Change
  • 276797 05/27/09 07:30 miham Beam current change
  • 276796 05/27/09 07:25 S. Frullani Neutron Detector N1 Right Channel 0
  • 276791 05/27/09 06:46 miham Target strip charts after current change
  • 276790 05/27/09 06:44 miham Spot++ check after Beam current change
  • 276789 05/27/09 06:37 R. Subedi Compton looks good
  • 276788 05/27/09 06:35 miham Beam current increased to 12uA
  • 276787 05/27/09 06:30 R. Subedi Scaler rates @12uA
  • 276786 05/27/09 06:30 miham Started New Compton Run #17035
  • 276780 05/27/09 05:52 miham Target Strip charts
  • 276776 05/27/09 05:49 miham Laser absorption spectrum 05-27-09 (05:51)
  • 276775 05/27/09 05:48 miham Spot++ check (#2807)
  • 276753 05/27/09 03:57 miham New Compton Run (#17034)
  • 276748 05/27/09 03:32 miham Target strip charts
  • 276747 05/27/09 03:29 miham Scaler snap shots (#2803)
  • 276746 05/27/09 03:27 miham Spot++ check (#2803)
  • 276741 05/27/09 03:06 s. Frullani Left arm Cherenkov 6
  • 276732 05/27/09 02:02 miham New Compton run #17033
  • 276719 05/27/09 00:44 miham Firefox problem on adaql[4-10] computers
  • 276718 05/27/09 00:37 miham Scaler snap shots (#2797)
  • 276717 05/27/09 00:33 miham Spot++ check (#2797)
  • 276716 05/27/09 00:30 miham ND, BB and HRS-R HV Gui screen shot - All Green
  • 276715 05/27/09 00:25 miham NMR frequency sweep (polarization 58.6% ) and Laser absorption spectrum.
  • 276710 05/27/09 00:19 miham Target Strip charts
  • 276709 05/27/09 00:15 miham New golden runs #2793 and #21693
  • 276708 05/27/09 00:02 tireman Swing shift summary
  • Tuesday

  • 276706 05/26/09 23:52 tireman Started new Compton Run
  • 276696 05/26/09 23:05 tireman Left HRS DAQ hiccups
  • 276695 05/26/09 23:00 tireman Harp Scan from beam recovery around 1800 hrs
  • 276694 05/26/09 22:54 X. Zheng Late entry - access activity at 9:30am and 12:30pm (PVDIS work)
  • 276689 05/26/09 22:05 tireman spot++ check after beam returned from controlled access
  • 276688 05/26/09 22:05 A. Camsonne Compton checks
  • 276687 05/26/09 22:03 tireman back to beam delivery
  • 276682 05/26/09 21:56 A. Camsonne Compton electron detector in
  • 276681 05/26/09 21:40 Bo/Ramesh RHRS scalar screen for cosmic run
  • 276680 05/26/09 21:30 tireman Data Runs since beam recovery: # 2787 - 2792 and # 21687 - 21692
  • 276679 05/26/09 21:16 A. Camsonne Compton cavity striptool for Sirish
  • 276676 05/26/09 21:08 R. Subedi Asked for controlles access
  • 276675 05/26/09 21:06 Yawei Right arm S1 TDC plots
  • 276672 05/26/09 20:44 R. Subedi RHRS scaler screen, T1 rate missing
  • 276671 05/26/09 20:41 R. Subedi T1 from RHRS missing
  • 276668 05/26/09 20:24 Abdurahim cavity parameter settings as of 05/26/09 afternoon
  • 276652 05/26/09 19:47 tireman Shift Checklist spot++ snapshot
  • 276651 05/26/09 19:44 tireman Shift Check list Hall A tool screen snapshot
  • 276650 05/26/09 19:40 tireman Run # 2788 and 21688 end early
  • 276649 05/26/09 19:39 Yawei Right S1 TDC pulser
  • 276644 05/26/09 19:28 Sirish Compton Cavity Stability
  • 276643 05/26/09 19:26 tireman HRS scaler screen snapshots
  • 276642 05/26/09 19:20 J. Watson Runs 2787 and 21687 ended early
  • 276639 05/26/09 19:14 Yawei Start new happex run , run#31425
  • 276636 05/26/09 18:49 Yawei NMR frequency sweep , polarization 58.13%. Absorption spectrum.
  • 276635 05/26/09 18:49 tireman Production running restarted
  • 276632 05/26/09 18:42 tireman Spot++ check for 5.5 micro amp beam and 4x4 raster
  • 276630 05/26/09 18:38 tireman Scaler check for beam deliever
  • 276628 05/26/09 18:31 tireman He-3 spot++ check with raster at 6x6
  • 276624 05/26/09 18:25 tireman He-3 target scaler check out
  • 276623 05/26/09 18:20 tireman BeO optics check is good
  • 276621 05/26/09 18:11 tireman Spot++ for beam delievery checkout
  • 276620 05/26/09 18:10 tireman Beam delievery checkout begins
  • 276619 05/26/09 18:09 Yawei Target moved to 3He
  • 276616 05/26/09 18:02 Yawei Target moved to BeO
  • 276615 05/26/09 17:55 Yawei Target moved to empty position
  • 276614 05/26/09 17:51 Yawei Target moved to BeO
  • 276613 05/26/09 17:26 ellie Changed HAND HV
  • 276612 05/26/09 17:26 ellie Changed HAND HV
  • 276611 05/26/09 17:08 Vince/Miha New T2 scaler copy put into the End-of-run script.
  • 276610 05/26/09 16:14 Yawei Target moved to empty position
  • 276605 05/26/09 16:08 grames Hall A QE still good
  • 276604 05/26/09 16:04 S. Sirca Day shift SUMMARY
  • 276600 05/26/09 14:38 L. Zhu target: screen shot of the laser absorption spectrum
  • 276599 05/26/09 14:34 L. Zhu target: screen shot of laser control windows
  • 276598 05/26/09 12:57 Jin Huang More EPR Calibration with Correction Coil Current Scan
  • 276597 05/26/09 12:52 S. Sirca Right HRS at 2.175 GeV/c (and at 18.0 deg).
  • 276596 05/26/09 12:17 Bo overcounting scalar check in RHRS
  • 276586 05/26/09 11:58 Jin Huang EPR Calibration for Transverse- Field of Cell Moss; Correct Coils are turned on
  • 276585 05/26/09 11:47 Higinbotham BigBite Field Off Briefly
  • 276584 05/26/09 11:36 Higinbotham BigBite Set To 82 Degrees
  • 276583 05/26/09 11:22 Higinbotham Right Arm 18 Degrees
  • 276580 05/26/09 09:58 gjin gas shed: argon/ethane one bottle empty
  • 276579 05/26/09 09:27 Moffit HAPPEX DAQ: restart vncviewer
  • 276578 05/26/09 09:11 Moffit HAPPEX DAQ: changes...
  • 276575 05/26/09 08:43 L. Zhu status: controlled access
  • 276568 05/26/09 08:06 miham OWL Shift Summary (05/26/09)
  • 276562 05/26/09 07:40 miham Target Strip charts
  • 276561 05/26/09 07:35 miham NMR frequency sweep (polarization 55.5% ) and Laser absorption spectrum.
  • 276528 05/26/09 04:11 miham Scaler snap shots for BB, HRS-L, and HRS-R
  • 276518 05/26/09 03:57 miham Spot++ check (#2769)
  • 276517 05/26/09 03:54 miham Compton is not running
  • 276512 05/26/09 03:19 miham Target Strip Charts
  • 276511 05/26/09 03:16 miham NMR Frequency Sweep , polarization 55.9%
  • 276487 05/26/09 00:46 miham Spot++ check (#2763)
  • 276485 05/26/09 00:43 miham Laser absorption spectrum 05-26-09 (00:26) - This time Correct picture
  • 276484 05/26/09 00:28 miham BB, ND, HRS-R HV screen - All Green
  • 276483 05/26/09 00:25 miham Laser absorption spectrum 05-26-09 (00:26)
  • 276482 05/26/09 00:22 miham Target Strip Charts.
  • 276477 05/26/09 00:01 tireman Swing shift summary
  • Monday

  • 276475 05/25/09 23:59 M.Shabestari Target--->Strip charts
  • 276474 05/25/09 23:57 M.Shabestari Laser absorption spectrum
  • 276469 05/25/09 23:21 M.Shabestari Laser absorption spectrum
  • 276462 05/25/09 22:52 tireman HAPPEX run started 31418
  • 276459 05/25/09 22:46 tireman Run # 2758 and 21662 no good ... no Left HRS events
  • 276456 05/25/09 22:24 M.Shabestari Target--->Strip charts
  • 276454 05/25/09 21:57 tireman Beam off for 25 minutes
  • 276451 05/25/09 21:48 tireman Run # 2755 and 21659 ended early due to lengthy beam off
  • 276450 05/25/09 21:44 tireman Run #'s 2757 and 21661 no events due to long beam loss
  • 276439 05/25/09 20:53 M.Shabestari Laser absorption spectrum
  • 276434 05/25/09 20:35 tireman Hall A tools snapshot
  • 276433 05/25/09 20:32 tireman spot++ screen shots
  • 276419 05/25/09 20:02 S. Sirca Wrong run status in shift summary
  • 276414 05/25/09 19:10 tireman HAPPEX run 31417 started
  • 276411 05/25/09 19:09 M.Shabestari NMR frequency sweep
  • 276401 05/25/09 18:09 Jin Huang Replay Scripts Update
  • 276399 05/25/09 18:07 M.Shabestari Target---> strip charts
  • 276394 05/25/09 17:33 R. Subedi Beam Time accounting problem
  • 276393 05/25/09 17:24 Yawei right arm VDC plots
  • 276392 05/25/09 17:17 Yawei left arm VDC plots
  • 276383 05/25/09 16:15 M.Shabestari Laser absorption spectrum
  • 276378 05/25/09 16:11 S. Sirca Day shift SUMMARY
  • 276372 05/25/09 15:19 S. Sirca Start new HAPPEX run #31416
  • 276371 05/25/09 15:19 brad sch runs 2743 &21647 are junk
  • 276364 05/25/09 15:07 brad sch runs 2742 & 21646 are junk
  • 276363 05/25/09 15:07 S. Sirca Started new HAPPEX run #31415
  • 276358 05/25/09 15:01 pkpan NMR 56%
  • 276317 05/25/09 11:23 S. Sirca Hall A General Tools
  • 276316 05/25/09 11:20 S. Sirca Spot++ check during run #2734
  • 276312 05/25/09 10:55 S. Sirca Started new HAPPEX run #31414
  • 276311 05/25/09 10:55 pkpan NMR_56%
  • 276296 05/25/09 08:52 V. Sulkosky R-HRS T1 [++] Over Counting
  • 276291 05/25/09 08:17 ellie Owl Shift Summary
  • 276285 05/25/09 07:41 ellie Started New Compton Run # 17024
  • 276283 05/25/09 07:40 ellie Started New HAPPEX Run # 31413
  • 276277 05/25/09 07:07 ellie MCC restored beam
  • 276274 05/25/09 06:55 ellie MCC called, down for a few minutes for an optics check
  • 276273 05/25/09 06:30 R. Michaels NMR 55.9 %
  • 276262 05/25/09 05:25 ellie Beam has returned!
  • 276261 05/25/09 05:14 ellie Beam Down due to MCC RF problems
  • 276258 05/25/09 05:09 R. Michaels absorption spectrum (again)
  • 276257 05/25/09 05:00 ellie Started new Compton Run # 17023
  • 276252 05/25/09 04:51 ellie VDC Efficiency Screen Grab
  • 276251 05/25/09 04:32 R. Michaels target is ok. beam back
  • 276250 05/25/09 04:30 ellie Beam Restored
  • 276247 05/25/09 04:26 ellie Beam off to fix Target
  • 276240 05/25/09 04:14 ellie Hall A Tools Screen Grab
  • 276239 05/25/09 04:11 ellie Spot++ Screen Grab
  • 276238 05/25/09 04:08 R. Michaels absorption spectrum
  • 276224 05/25/09 03:24 R. Michaels NMR 56 %
  • 276217 05/25/09 02:51 ellie New HAPPEX Run # 31412
  • 276214 05/25/09 02:28 ellie Compton End of Run failed, rebooted, Started Run # 17022
  • 276212 05/25/09 02:24 ellie ROC8 stopped responding, restarted crate, back up and running
  • 276202 05/25/09 01:25 miham Scaler comparison
  • 276193 05/25/09 00:31 ellie Restarted Compton, few fails, back on with Run # 17021
  • 276188 05/25/09 00:01 tireman Swing shift summary
  • Sunday

  • 276182 05/24/09 23:28 Yawei NMR frequency sweep , polarization 55.9%. Absorption spectrum.
  • 276177 05/24/09 23:00 tireman Started HAPPEX run 31411
  • 276160 05/24/09 21:21 tireman tools screen snapshot
  • 276159 05/24/09 21:19 miham Target Polarization during (e',d)
  • 276158 05/24/09 21:17 tireman spot++ for checklist
  • 276132 05/24/09 19:14 Jin Huang Spin Signal Changed from "-" to "+" before production 2701/21605
  • 276131 05/24/09 19:13 Jin Huang Spin Signal Changed from "-" to "+" before production 2701/21605
  • 276126 05/24/09 18:51 Jin Huang Spin Rotated to Transverse+ (90 degree) before production 2270/21604
  • 276125 05/24/09 18:42 tireman Happex run # 31409 stopped #31410 started
  • 276124 05/24/09 18:39 tireman Target sping flip
  • 276119 05/24/09 18:24 Yawei NMR frequency sweep , polarization 58.17%. Absorption spectrum.
  • 276118 05/24/09 18:20 Yawei NMR frequency sweep , polarization 58.17%. Absorption spectrum.
  • 276101 05/24/09 16:24 tireman Started HAPPEX run # 31409
  • 276095 05/24/09 15:52 puckett day shift summary
  • 276082 05/24/09 14:33 puckett tools snapshot for checklist
  • 276081 05/24/09 14:32 A. Puckett spot++ for checklist
  • 276076 05/24/09 14:12 puckett started Compton run 17018 at 14:10
  • 276075 05/24/09 13:41 St. John NMR frequency sweep , polarization 57.8%. Absorption spectrum.
  • 276062 05/24/09 12:17 puckett Ended HAPPEX run 31407, started HAPPEX run 31408 @ 12:04
  • 276030 05/24/09 09:34 St. John Last NMR Run
  • 276027 05/24/09 09:21 R. Subedi Target Polarization Table
  • 276026 05/24/09 09:20 St. John NMR frequency sweep , polarization 58.4%. Absorption spectrum.
  • 276017 05/24/09 08:26 puckett started HAPPEX run 31407 at 8:06
  • 276016 05/24/09 08:25 puckett started Compton run 17017
  • 276010 05/24/09 08:01 ellie Owl Shift Summary
  • 275993 05/24/09 06:11 ellie Compton restarted, Run # 17016
  • 275992 05/24/09 05:49 ellie DAQs Rebooted, some junk runs occured
  • 275986 05/24/09 05:42 Nelyubin NMR frequency sweep , polarization 57.7%. Absorption spectrum.
  • 275971 05/24/09 04:14 ellie Started new Compton run # 17015
  • 275968 05/24/09 04:13 ellie Started new HAPPEX run # 31406
  • 275965 05/24/09 04:10 ellie VDC Efficiency Screen Grabs
  • 275955 05/24/09 03:55 ellie Hall A Tools Screen Grab
  • 275954 05/24/09 03:54 ellie Spot++ Screen Grabs
  • 275937 05/24/09 01:58 ellie Started new Compton run after a few failed attempts and CODA/ROC reboots
  • 275936 05/24/09 01:51 Nelyubin NMR Frequency Sweep , polarization 58.16%
  • 275933 05/24/09 01:43 ellie HAPPEX Restarted, current run # 31405
  • 275930 05/24/09 01:39 ellie Alarm Handler occassionally beeping once, then stopping with no errors shown
  • 275925 05/24/09 00:53 A. Camsonne Compton cavity still not locking
  • 275923 05/24/09 00:47 ellie RHRS DAQ Stopped Counting, Restarted CODA, everything looks fine now
  • Saturday

  • 275911 05/23/09 23:58 miham Compton Not working
  • 275910 05/23/09 23:56 puckett swing shift summary
  • 275909 05/23/09 23:56 puckett no--Compton not working.
  • 275900 05/23/09 23:03 M.Shabestari Target--->Strip charts
  • 275899 05/23/09 23:02 M.Shabestari Laser absorption spectrum
  • 275894 05/23/09 22:44 puckett belated entry: HAPPEX run 31403 ended/run 31404 started at 21:48
  • 275887 05/23/09 21:51 M.Shabestari NMR Frequency Sweep
  • 275880 05/23/09 21:33 puckett spot++ for checklist
  • 275879 05/23/09 21:24 miham BB and HRS-L Scaler snap shots
  • 275870 05/23/09 20:40 M.Shabestari Target---> Strip charts
  • 275867 05/23/09 20:18 M.Shabestari Laser absorption spectrum
  • 275866 05/23/09 20:18 brads Is the Compton even working?
  • 275842 05/23/09 18:47 puckett new Compton run 17011--started by Miha
  • 275837 05/23/09 18:22 puckett This might be my fault--recent event on adaql8
  • 275834 05/23/09 17:41 M.Shabestari Target---> Strip charts
  • 275833 05/23/09 17:35 M.Shabestari NMR Frequency Sweep
  • 275832 05/23/09 17:32 puckett Ended HAPPEX run 31402, started HAPPEX run 31403 @17:31
  • 275831 05/23/09 17:30 brads Restarted auto-snapshot software
  • 275819 05/23/09 17:03 M.Shabestari Laser Absoption Spectrum
  • 275809 05/23/09 16:05 X. Zheng Day shift summary
  • 275792 05/23/09 14:16 X. Zheng spot++ run 2639
  • 275787 05/23/09 13:57 R. Subedi Target polarization increased!
  • 275786 05/23/09 13:34 X. Zheng start HAPPEX run 31402
  • 275783 05/23/09 13:29 A. Deur NMR
  • 275782 05/23/09 13:27 X. Zheng Insert beam HWP
  • 275781 05/23/09 13:22 X. Zheng stopped HAPPEX run 31401
  • 275778 05/23/09 13:18 A. Deur Target temperatures, fields and polarization screenshot
  • 275777 05/23/09 13:15 A. Deur Laser absorption spectrum
  • 275748 05/23/09 09:53 tireman Trigger 1 Left-HRS prescale incorrectly set over night (5-22-09 thru 5-23-09)
  • 275747 05/23/09 09:41 tireman Runs 2629 and 21532 junk
  • 275744 05/23/09 09:33 X. Zheng start HAPPEX run 31401
  • 275734 05/23/09 08:52 miham LabView restart
  • 275733 05/23/09 08:37 A. Deur NMR done after labiew restart (8:31am May 23rd)
  • 275730 05/23/09 08:30 dwang NMR
  • 275725 05/23/09 07:59 Ibrahim Owl Shift Summary (May 23, 2009)
  • 275717 05/23/09 07:06 Yi Qiang Labview programs frozen
  • 275712 05/23/09 06:54 Nelyubin NMR 51.64% 2200_05_23 6:28
  • 275703 05/23/09 05:34 Ibrahim Tools Screen for the Checklist
  • 275698 05/23/09 05:18 Ibrahim Spot++ for the Checklist
  • 275697 05/23/09 05:16 Ibrahim Started the HAPPEX DAQ. New Run 31400
  • 275690 05/23/09 04:47 Nelyubin 3He polarized target on the beam
  • 275685 05/23/09 04:13 Nelyubin H2 reference target on beam at 134 psig
  • 275676 05/23/09 03:52 Nelyubin Optics target on beam
  • 275671 05/23/09 03:07 Higinbotham Optics Data During Spin-Up
  • 275666 05/23/09 02:46 Nelyubin H2 reference target on beam
  • 275661 05/23/09 02:29 Nelyubin Optics target on beam
  • 275660 05/23/09 02:26 Nelyubin NMR sweep , polarization 38.5%
  • 275659 05/23/09 02:23 miham Target moved from 3He to Optics
  • 275658 05/23/09 02:12 Nelyubin set 3He cell on beam
  • 275653 05/23/09 01:58 Nelyubin Set optics target on beam
  • 275642 05/23/09 00:47 Nelyubin H2 reference cell at 134 PSI on beam
  • 275641 05/23/09 00:46 Nelyubin H2 reference cell at 134 PSI on beam
  • 275631 05/23/09 00:02 A. Puckett swing shift summary
  • Friday

  • 275630 05/22/09 23:44 puckett tools snapshot
  • 275629 05/22/09 23:35 puckett Moved to BeO--taking data at 2 uA for now, waiting for instructions
  • 275624 05/22/09 23:20 brad sch rate check @ 4x4 raster
  • 275623 05/22/09 23:18 puckett spot++ at 4x4 raster size @ 2 uA
  • 275618 05/22/09 23:13 brad sch rate check @ 2x2 raster
  • 275617 05/22/09 23:10 puckett spot++ @ 2x2 raster size
  • 275612 05/22/09 23:03 brad sch rate check @ 6x6 raster
  • 275611 05/22/09 23:00 puckett spot++ @ 6x6 raster
  • 275606 05/22/09 22:44 Yi Qiang First NMR after the target access: 16.7%
  • 275605 05/22/09 22:42 Yi Qiang First NMR after the target access: 19%
  • 275604 05/22/09 22:41 adaq Trig. rate snapshots, 2 uA, standard raster size
  • 275601 05/22/09 22:29 Yi Qiang Will keep target pumping as it is, no more changes tonight
  • 275600 05/22/09 22:26 Jin Huang 1/4 plates calibration for Transverse Laser
  • 275599 05/22/09 22:25 Yi/Jin Laser power test
  • 275598 05/22/09 22:05 Yi/Jin Target activities in the hall
  • 275595 05/22/09 21:01 miham New copy of BB scalers added to HRS-L scalers
  • 275594 05/22/09 20:56 ellie Changed HAND V-22L HV
  • 275593 05/22/09 20:51 brad sch runs 2601 & 21504 are cosmic
  • 275586 05/22/09 18:30 dwang NMR
  • 275564 05/22/09 16:03 X. Zheng day shift summary
  • 275561 05/22/09 15:54 X. Zheng HAPPEX run 31399 started at 3:54pm
  • 275556 05/22/09 15:08 R. Michaels absorption spectrum
  • 275539 05/22/09 13:14 X. Zheng spot++ run 2590
  • 275538 05/22/09 13:03 R. Michaels NMR 52.28
  • 275525 05/22/09 11:50 tireman Run 21489 Junk
  • 275524 05/22/09 11:49 X. Zheng start HAPPEX new run 31398
  • 275521 05/22/09 11:35 R. Michaels added HWP to HAPPEX data
  • 275508 05/22/09 09:55 X. Zheng beam away
  • 275507 05/22/09 09:18 R. Michaels NMR 51.588
  • 275502 05/22/09 08:46 tireman Run 21484 end early
  • 275501 05/22/09 08:45 tireman Run 2582 end early
  • 275500 05/22/09 08:45 R. Michaels L-HRS DAQ died again
  • 275496 05/22/09 08:23 R. Michaels DAQ reboot
  • 275488 05/22/09 07:58 X. Zheng start HAPPEX run 31397
  • 275487 05/22/09 07:53 Nelyubin Laser absorption spectrum 05/22/2009 7:50
  • 275486 05/22/09 07:52 ellie Owl Shift Summary
  • 275463 05/22/09 05:06 Nelyubin NMR Frequency Sweep 51.82%
  • 275460 05/22/09 04:45 ellie VDC Efficiency Screen Grabs
  • 275457 05/22/09 04:28 ellie Hall A General Tools Screen Grab
  • 275454 05/22/09 04:27 ellie Spot++ Screen Grabs
  • 275444 05/22/09 04:08 ellie Restarted HAPPEX, Current Run # 31396
  • 275415 05/22/09 01:04 Nelyubin NMR Frequency Sweep 51.53%
  • 275414 05/22/09 00:56 ellie Typo, Previous post should read "BCM"
  • 275410 05/22/09 00:50 ellie BPM looks like it's drifting
  • 275403 05/22/09 00:17 Norum Shift summary
  • Thursday

  • 275398 05/21/09 23:52 gjin trigger type distrubution: changed?
  • 275397 05/21/09 23:37 R. Subedi ROC5 error message
  • 275395 05/21/09 23:30 miham HRS-L & BB Scaler snap shots
  • 275391 05/21/09 23:04 Norum DAQ Problem
  • 275372 05/21/09 21:11 dwang NMR Sweep
  • 275348 05/21/09 19:24 miham Begin-of-run files missing
  • 275345 05/21/09 19:17 miham Spot++ Check on 3He before production
  • 275338 05/21/09 18:55 miham Spot++ check on 3He, (6x6)mm
  • 275333 05/21/09 18:50 miham Spot++ check on 3He , (2x2)mm
  • 275330 05/21/09 18:43 miham Spot++ check on BeO/Optics, 2uA, 4x4mm
  • 275321 05/21/09 18:26 A. Camsonne Compton cavity not locking
  • 275320 05/21/09 18:13 miham ND HV trip
  • 275317 05/21/09 17:17 miham Target Strip Charts before beam
  • 275316 05/21/09 17:13 Brad & Miha Missing Start-of-run file for run #21449 - Cosmics run
  • 275307 05/21/09 17:01 miham NMR measurement before Beam (pol. is 47.7%)
  • 275304 05/21/09 15:53 A. Camsonne Compton electron detector in garage position
  • 275303 05/21/09 15:34 A. Camsonne Compton electron detector in beam position for test
  • 275302 05/21/09 15:23 miham JUNK runs!!!
  • 275300 05/21/09 15:07 Anderson NMR Measurement 43.9%
  • 275299 05/21/09 15:07 dezern Checklist
  • 275298 05/21/09 14:48 R. Michaels RHRS EDTM timing check
  • 275297 05/21/09 12:56 Yi Qiang First NMR after correction coil turned ON: 29%
  • 275296 05/21/09 12:56 Yi Qiang EPR snap shots
  • 275295 05/21/09 12:54 Yi Qiang EPR Calibration for Transverse+ field with Correction coil ON
  • 275294 05/21/09 12:44 Yi Qiang AFP Loss test for Transverse+ field
  • 275283 05/21/09 11:57 ellie Changed HAND V-22L HV
  • 275282 05/21/09 11:03 Yi/Jin Transverse laser alignment checked
  • 275281 05/21/09 10:32 Bo Wide pedestal of Preshower ADC44 in RHRS
  • 275272 05/21/09 09:47 V. Sulkosky Adjusted Lumi HV's
  • 275271 05/21/09 09:45 X. Zheng access activity [PVDIS work]
  • 275270 05/21/09 09:43 Doug & Miha Checked V22-L Signal
  • 275267 05/21/09 09:33 V. Sulkosky Pedestal Runs for Left and Right HRS
  • 275266 05/21/09 09:06 Higinbotham Logging Which Transverse Direction
  • 275261 05/21/09 08:26 R. Subedi Hall in restricted access, please take access only after coordination with either SL or RC
  • 275259 05/21/09 08:02 jpchen owl shift summary
  • 275258 05/21/09 07:16 jpchen tool screen
  • 275257 05/21/09 07:15 R. Subedi Beam study starts, asked for controlled access
  • 275256 05/21/09 07:07 D. Parno NMR measurement: 51.2%
  • 275255 05/21/09 07:01 jpchen Beam off for RF recovery
  • 275252 05/21/09 06:51 R. Subedi BLumi5 still saturating [lumi experts please look]
  • 275239 05/21/09 05:14 D. Parno Compton restart: 16993. Locking better in last 30 minutes.
  • 275238 05/21/09 05:13 jpchen started happex run 31394
  • 275216 05/21/09 04:01 D. Parno Laser absorption spectrum
  • 275215 05/21/09 03:51 D. Parno New Compton run 16992
  • 275196 05/21/09 01:45 D. Parno NMR Measurement: 50.1%
  • 275193 05/21/09 01:42 jpchen spot check 6x6 on pol. He3
  • 275188 05/21/09 01:38 jpchen spot check 4x4 on pol. He3
  • 275187 05/21/09 01:36 D. Parno Target was moved to polarized 3He at 01:30
  • 275184 05/21/09 01:34 jpchen start happex run 31392
  • 275183 05/21/09 01:33 jpchen spot check raster 4x4 on BeO
  • 275181 05/21/09 01:32 jpchen harp scan
  • 275180 05/21/09 01:31 D. Parno New Compton run 16991
  • 275175 05/21/09 01:18 jpchen Beam checking: spot with raster off on BeO
  • 275170 05/21/09 00:30 D. Parno Target moved to BeO
  • 275168 05/21/09 00:01 pking Swing shift summary
  • Wednesday

  • 275167 05/20/09 23:11 dwang NMR May 20, 23:10
  • 275164 05/20/09 23:06 pking Communication with iochla not restored.
  • 275163 05/20/09 22:51 pking OPS lost communication with iochla.
  • 275158 05/20/09 21:49 D. Parno Compton restart; new run 16990
  • 275157 05/20/09 21:44 miham New Thresholds for BB MWDC (5.5V) - Back to production
  • 275150 05/20/09 21:39 miham New thresholds for the BB MWDC (4.0V)
  • 275146 05/20/09 21:34 miham New Thresholds for BB MWDC (4.25V)
  • 275141 05/20/09 21:29 miham New Thresholds for BB MWDC (4.75V)
  • 275136 05/20/09 21:24 miham New thresholds for BB MWDC (5.0V)
  • 275131 05/20/09 21:19 miham New Thresholds for BB MWDC (5.25V)
  • 275127 05/20/09 21:09 pking Hall A general tools
  • 275126 05/20/09 21:00 pking Spot++ check, raster shows 4x4
  • 275108 05/20/09 19:50 pking BLumi5 versus beam current during a trip.
  • 275103 05/20/09 19:45 pking Started new Compton run, 16989, at 19:44.
  • 275102 05/20/09 19:43 pking BLumi5 seems to be saturating.
  • 275101 05/20/09 19:40 pking Started new HAPPEX run, 31391, at 19:39
  • 275096 05/20/09 19:07 Yi Qiang NMR Before and After Spin Rotation: 54.5 and 43.5%
  • 275095 05/20/09 19:05 Yi/Jin EPR Calibration for Transverse+ Field of Cell Moss (with correct snapshots)
  • 275094 05/20/09 19:05 Yi/Jin EPR Calibration for Transverse+ Field of Cell Moss
  • 275093 05/20/09 18:58 Yi/Jin Spin Rotated to Transverse- (270 degree) before production 2514/21418
  • 275084 05/20/09 17:44 pking Started new Compton run, 16988, at 17:42
  • 275071 05/20/09 16:09 R. Subedi RHRS DAQ running now, thanks Bryan
  • 275067 05/20/09 16:03 X. Zheng Shift Summary
  • 275066 05/20/09 15:46 R. Subedi RHRS DAQ still problem, running LHRS only now
  • 275064 05/20/09 15:32 R.Subedi RHRS DAQ problem
  • 275063 05/20/09 15:19 X. Zheng spot++ with 6x6 raster
  • 275059 05/20/09 15:11 Anderson NMR Measurement 53.6% Transverse -
  • 275058 05/20/09 15:05 X. Zheng spot++ on BeO
  • 275055 05/20/09 13:38 A. Saha Moller Measurement
  • 275054 05/20/09 13:34 X. Zheng BB Wire Chamber HV
  • 275053 05/20/09 13:32 X. Zheng Starting beam recovery procedure
  • 275052 05/20/09 12:21 miham MWDC Threshold scan results
  • 275042 05/20/09 10:42 Anderson NMR Measurement 53.3% Transverse -
  • 275041 05/20/09 10:27 X. Zheng Prepare for Moller (BBWC HV off, empty target)
  • 275030 05/20/09 08:52 riordan LHRS/BigBite Golden Run 2428
  • 275025 05/20/09 08:25 Jack (Techs) Right Arm Q1 power supply
  • 275024 05/20/09 08:10 V. Sulkosky Re: RHRS Preshower ADC 44
  • 275023 05/20/09 08:02 ellie Owl Shift Summary
  • 275017 05/20/09 07:37 ejensen ROC Sync Check
  • 275010 05/20/09 06:57 D. Parno NMR Measurement: 51.8%
  • 275005 05/20/09 06:25 ellie Compton restarted, new run: 16985
  • 274990 05/20/09 04:46 ellie Hall A Tools Screen Grab
  • 274989 05/20/09 04:43 ellie Spot++ Screen Grab
  • 274986 05/20/09 04:28 ellie Restarted Compton (16984) and HAPPEX (31388)
  • 274974 05/20/09 04:13 D. Parno Laser absorption spectrum
  • 274973 05/20/09 04:04 ellie VDC Efficiencies Screen Grabs
  • 274968 05/20/09 03:46 ejensen Event type distribution
  • 274967 05/20/09 03:41 ellie "Mystery" (aka Compton Photon Detector) HV Tripped
  • 274966 05/20/09 03:28 ellie Previous entry should read "off and on", not the other way around
  • 274961 05/20/09 03:20 ellie Turned HV for V-22L on and off in between runs
  • 274960 05/20/09 03:02 ellie LHRS Sync Gaps for Run 2488
  • 274959 05/20/09 02:56 D. Parno NMR Measurement: 49.9%
  • 274949 05/20/09 02:00 ellie Compton Restarted, Current Run: 16983
  • 274944 05/20/09 01:54 ellie HAND N1 Left TDCs Working, disregard HALog 274932
  • 274943 05/20/09 01:26 ellie Changed RHRS Golden Run to 21393
  • 274942 05/20/09 01:26 ellie What's up with RHRS Preshower ADC 44?
  • 274937 05/20/09 01:01 ellie HAND V-22L Problems Return
  • 274932 05/20/09 00:36 ellie Plane 1 Left TDCs Not Working - Bad Discriminator?
  • 274931 05/20/09 00:33 Franklin Compton
  • 274928 05/20/09 00:09 ellie Restarted HAPPEX (current run #31387) and Compton (Current run #16982)
  • 274924 05/20/09 00:05 jpchen swing shift summary
  • Tuesday

  • 274919 05/19/09 23:38 NMR polarization measured before run 2483 and 21392: 46.7%
  • 274918 05/19/09 23:18 Yi Qiang EPR Calibration for Transverse- Field of Cell Moss
  • 274913 05/19/09 22:38 Moffit LeftHRS T2 in the helicity/target gated scalers.
  • 274912 05/19/09 22:37 jpchen Polarization lower than expected
  • 274911 05/19/09 22:33 Parno/Frankl Original Compton Analysis Failing for Prod_FA and Prod_FT
  • 274904 05/19/09 22:02 Kai Pan NMR 51% 2200_5_19
  • 274903 05/19/09 21:58 Kai Pan Target moves to the He3
  • 274894 05/19/09 21:29 Kai Pan Target moves to Optics
  • 274891 05/19/09 21:21 D. Parno Now running Compton in normal production (Prod_FA mode)
  • 274890 05/19/09 20:58 D. Parno Stripchart figure for Entry 274889
  • 274889 05/19/09 20:54 D. Parno Compton running again
  • 274875 05/19/09 20:06 B. Norum Checkout, Raster on
  • 274870 05/19/09 19:54 jpchen raster 2x2 on reference H2
  • 274865 05/19/09 19:48 jpchen raster 6x6 on reference H2
  • 274860 05/19/09 19:40 jpchen raster 4x4 on H2
  • 274857 05/19/09 19:36 Kai Pan Target moves to the reference cell
  • 274856 05/19/09 19:35 Kai Pan Filled 135psig H2 into the reference cell
  • 274855 05/19/09 19:32 jpchen raster check 4x4 BeO
  • 274850 05/19/09 19:28 jpchen harp scan results
  • 274847 05/19/09 19:16 R. Michaels unswap DELAY3 and 4 (late entry)
  • 274846 05/19/09 19:16 jpchen spot check with raster off
  • 274845 05/19/09 19:13 jpchen Beam back
  • 274839 05/19/09 17:39 Kai Pan Target moves to He3
  • 274838 05/19/09 17:38 atrig Trigger-Download
  • 274836 05/19/09 17:33 jpchen Need to replace Turbo: estimation two hours down
  • 274834 05/19/09 17:30 atrig Trigger-Download
  • 274832 05/19/09 17:29 atrig Trigger-Download
  • 274831 05/19/09 17:12 jpchen NL access to restart turbo pump at 1L06
  • 274828 05/19/09 16:51 R. Michaels swap DELAY3, DELAY4
  • 274825 05/19/09 16:37 jpchen Beam down for ~ 1hour for NL access
  • 274824 05/19/09 16:36 jpchen raster check
  • 274821 05/19/09 16:29 jpchen Harp scan
  • 274816 05/19/09 16:25 jpchen Beam back, position checked
  • 274814 05/19/09 16:10 Higinbotham Day Shift Summary
  • 274813 05/19/09 16:04 Higinbotham Late Entry: BCM Calibration
  • 274812 05/19/09 15:55 Moffit / mih T2 double pulsing problem solved
  • 274811 05/19/09 15:55 Yi Qiang AFP loss of Transverse- field with corrention coil off
  • 274810 05/19/09 15:47 Yi Qiang Current target polarization: 43%
  • 274809 05/19/09 15:43 Yi Qiang Very late entry: AFP loss of longitudinal- field with low polarization and corrention coil off
  • 274806 05/19/09 15:32 A. Camsonne Compton electron detector to garage position
  • 274805 05/19/09 15:31 A. Camsonne Compton electron detector to garage position
  • 274804 05/19/09 15:16 hyde Compton Polarimeter eDet
  • 274803 05/19/09 15:14 Yi Qiang H2 bottle connected to the reference cell (D2 bottle disconnected)
  • 274802 05/19/09 15:13 Moffit HAND EDTM shift
  • 274791 05/19/09 14:43 R. Subedi Re: Access activity [PVDIS work]
  • 274790 05/19/09 14:43 A. Camsonne ETROC firmware loaded to latest version for check with oscilloscope
  • 274789 05/19/09 14:40 ellie HAND V-22L ADC/TDC Problems Revisited
  • 274788 05/19/09 14:32 Anderson NMR Measurement 36.9% Transverse -
  • 274787 05/19/09 14:31 ellie HAND N1-13L EDTM Signal/TDC Problems
  • 274779 05/19/09 13:29 Yi Qiang EPR Calibration for Transverse- Field of Cell Moss
  • 274778 05/19/09 13:00 Yi Qiang 3He spin rotated to transverse+ (90 degree)
  • 274777 05/19/09 12:43 Higinbotham Strange Helicity Information
  • 274776 05/19/09 12:40 Yi Qiang Internal Temperature Test
  • 274775 05/19/09 12:39 Yi Qiang Reference Cell Vented
  • 274774 05/19/09 12:33 Franklin FADC VME crate modification
  • 274757 05/19/09 12:06 R. Subedi Access activity [PVDIS work]
  • 274753 05/19/09 11:56 R. Michaels possible reason for EDTM problem
  • 274752 05/19/09 11:53 ellie HAND V-22L PMT HV Back On
  • 274749 05/19/09 11:51 ellie Turning off HAND V-22L for troubleshooting
  • 274746 05/19/09 11:42 atrig Trigger-Download
  • 274745 05/19/09 11:40 R. Michaels EDTM: attempt swap, continued wierdness
  • 274732 05/19/09 11:06 Yi Qiang NMR Cross Calibration of Longitudinal
  • 274727 05/19/09 11:03
  • 274722 05/19/09 10:35 Yi Qiang EPR Calibration for Longitudinal Pumping of Cell Moss
  • 274721 05/19/09 10:33 Moffit HAPPEX DAQ: Run Numbers stopped incrementing since 05/17/09 7:18am
  • 274720 05/19/09 10:31 Moffit Lumi's looked very strange for start of 31385
  • 274719 05/19/09 10:04 Anderson Happex run ended
  • 274718 05/19/09 09:45 Yi Qiang NMR Frequency Sweep: 61.97%
  • 274717 05/19/09 08:14 A. Tobias Owl Shift Summary: Tuesday 19 May 2009
  • 274715 05/19/09 06:58 A. Tobias Beam away for studies - Controlled Access requested for Hall A
  • 274708 05/19/09 05:58 ejensen ROC sync check
  • 274706 05/19/09 05:47 A. Tobias HAPPEX restarted: current run = 31396 started at 05:45
  • 274705 05/19/09 05:46 D. Parno NMR Measurement: 61.2%
  • 274701 05/19/09 05:39 A. Tobias hapc7 system time ahead 32 min
  • 274700 05/19/09 05:34 A. Tobias Beam Current History last 80 min
  • 274699 05/19/09 05:33 A. Tobias Hall A Tools and Right Q1 Momentum History
  • 274677 05/19/09 03:56 A. Tobias Spot++ Run 2449 w/ 4x4mm & 10uA
  • 274676 05/19/09 03:55 D. Parno Laser absorption spectrum
  • 274675 05/19/09 03:54 A. Tobias Right Dipole LHe Level
  • 274670 05/19/09 03:37 A. Tobias HVGUI all green
  • 274661 05/19/09 02:35 D. Parno Compton Cavity Performance Depends on Laser Polarization
  • 274656 05/19/09 01:58 A. Tobias Hall A Tools Snapshot & Right Q1 Momentum History
  • 274655 05/19/09 01:52 A. Tobias HAPPEX DAQ restarted: Current run = 31395 started at 01:49
  • 274653 05/19/09 01:51 D. Parno NMR Measurement: 61.5%
  • 274649 05/19/09 01:30 A. Tobias TS0 stopped responding - reboot TS0 & startcoda both HRS DAQs
  • 274635 05/19/09 00:24 D. Parno Compton update
  • 274630 05/19/09 00:02 miham SWING Shift Summary 05/18/09
  • Monday

  • 274624 05/18/09 23:31 M.Shabestari Target--->Strip charts
  • 274623 05/18/09 23:18 miham BB scaler snap shots #2439
  • 274622 05/18/09 23:18 miham Scaler snap shots during run #2439
  • 274621 05/18/09 23:15 miham Spot++ check during production run #2439
  • 274616 05/18/09 22:49 M.Shabestari Laser absorption spectrum
  • 274611 05/18/09 21:51 M.Shabestari NMR Frequency Sweep
  • 274608 05/18/09 21:45 miham BAD beam
  • 274605 05/18/09 21:30 miham HV GUI snap shot - All Green
  • 274604 05/18/09 21:12 miham Spot++ check during production run #2435
  • 274599 05/18/09 20:45 M.Shabestari Laser absorption spectrum 05-18-09 (20:43)
  • 274585 05/18/09 20:11 miham Dead time (DATAMON) is 6% / 16%
  • 274580 05/18/09 20:02 miham Momentum distribution in HRS-L before and after momentum change.
  • 274575 05/18/09 19:21 M.Shabestari Target---> Strip charts
  • 274570 05/18/09 18:47 M.Shabestari Laser absorption spectrum
  • 274561 05/18/09 18:06 miham Spot++ check during production run #2429
  • 274556 05/18/09 17:35 Higinbotham Left HRS Prescale
  • 274555 05/18/09 17:24 M.Shabestari Target--->Strip charts
  • 274554 05/18/09 17:22 M.Shabestari NMR Frequency Sweep
  • 274549 05/18/09 17:03 Sirish Compton Cavity Power
  • 274548 05/18/09 16:54 miham Scaler snap shots during run 2427
  • 274547 05/18/09 16:53 miham Spot++ check during production run 2427
  • 274542 05/18/09 16:47 M. Shabestar Laser absorption spectrum
  • 274541 05/18/09 16:26 A. Camsonne Reverted Compton Electron detector firmware to version trigger_ctrl_restored
  • 274535 05/18/09 16:01 jpchen day shift summary
  • 274530 05/18/09 15:41 jpchen Tools screen
  • 274529 05/18/09 15:40 jpchen spot++
  • 274510 05/18/09 13:24 jpchen Started Happex run 31392
  • 274509 05/18/09 13:23 jpchen Beam 1/2-wave plate out.
  • 274508 05/18/09 13:21 jpchen ended happex run 31391
  • 274505 05/18/09 13:11 Anderson Target Strip Charts
  • 274504 05/18/09 13:07 Anderson Laser absorption spectrum
  • 274503 05/18/09 12:52 Anderson NMR Measurement after D2 and Optics Runs 63.88%
  • 274500 05/18/09 12:48 jpchen moved target to pol. He3
  • 274495 05/18/09 12:12 jpchen moved to optics target
  • 274483 05/18/09 11:34 jpchen Ended Compton run 16959
  • 274478 05/18/09 11:01 jpchen time should be 10:50
  • 274477 05/18/09 10:56 jpchen 11:50 filled reference cell to 35 psig N2
  • 274474 05/18/09 10:51 Yi Qiang Temperature Change in the Hall
  • 274469 05/18/09 10:45 jpchen 10:00 evacuated reference cell
  • 274466 05/18/09 10:36 R. Michaels Run 2414 = empty target cell
  • 274457 05/18/09 09:06 jpchen Plan for D2, Empty, N2 and multi-12C runs.
  • 274456 05/18/09 09:03 jpchen stop/start happex/compton at ~8:25
  • 274455 05/18/09 08:43 jpchen reference cell re-filled 100 psig D2, re-adjusted perssure gauge
  • 274452 05/18/09 08:31 Anderson NMR Measurement 61.5%
  • 274448 05/18/09 08:00 A. Tobias Owl Shift Summary: Monday 18 May 2009
  • 274439 05/18/09 07:05 A. Tobias BPM positions and current snapshot
  • 274438 05/18/09 06:49 A. Tobias Hall A Tools Snapshot
  • 274437 05/18/09 06:47 A. Tobias Spot++ Run 2409 w/ 4x4mm raster & 10uA
  • 274432 05/18/09 06:23 A. Tobias Compton DAQ restarted: current run = 16958 started at 06:20
  • 274427 05/18/09 06:00 A. Tobias HRS-Right Q1 Momentum Snapshot
  • 274422 05/18/09 05:41 A. Tobias BPM positions snapshot
  • 274417 05/18/09 05:08 A. Tobias Hall A Tools Snapshot
  • 274412 05/18/09 04:28 A. Tobias Compton DAQ restarted: current run = 16957 started at 04:21
  • 274411 05/18/09 04:26 A. Tobias HAPPEX DAQ restarted: current run = 31390 started at 04:20
  • 274409 05/18/09 04:22 D. Parno Laser absorption spectrum
  • 274407 05/18/09 04:20 D. Parno NMR Measurement: 61.6%
  • 274387 05/18/09 03:07 A. Tobias Spot++ Run 2401 4x4mm 10uA
  • 274386 05/18/09 03:01 A. Tobias HRS-R stopped counting (run21304): reboot ROCs and coda
  • 274382 05/18/09 02:47 A. Tobias Hall A Tools Screenshot
  • 274376 05/18/09 02:30 A. Tobias Compton DAQ restarted: run = 16956
  • 274373 05/18/09 02:17 miham New thresholds for BB MWDC (5.5V) - Back to Production
  • 274372 05/18/09 02:15 adev BigBite WireChamber Threshold change!
  • 274367 05/18/09 02:09 miham New Thresholds for BB MWDC (5.75V)
  • 274364 05/18/09 02:04 adev BigBite WireChamber Threshold change!
  • 274361 05/18/09 02:02 miham New thresholds for BB MWDC (6.25V)
  • 274356 05/18/09 01:56 miham New Thresholds for BB MWDC (6.75V)
  • 274353 05/18/09 01:54 adev BigBite WireChamber Threshold change!
  • 274350 05/18/09 01:50 miham New Thresholds for BB MWDC (10.0V)
  • 274345 05/18/09 01:45 miham New Thresholds for BB MWDC (9.5V)
  • 274340 05/18/09 01:39 miham New Thresholds for BB MWDC (9.0V)
  • 274335 05/18/09 01:34 miham New Thresholds for BB MWDC (8.5V)
  • 274334 05/18/09 01:33 adev BigBite WireChamber Threshold change!
  • 274329 05/18/09 01:29 miham New Thresholds for BB MWDC (8.0V)
  • 274324 05/18/09 01:23 miham New Thresholds for BB MWDC (7.5V)
  • 274323 05/18/09 01:23 adev BigBite WireChamber Threshold change!
  • 274314 05/18/09 00:25 D. Parno NMR Measurement: 61.8%
  • 274311 05/18/09 00:09 LeRose Shift Summary (swing)
  • 274310 05/18/09 00:07 LeRose Q1 Current readback
  • Sunday

  • 274307 05/17/09 23:56 M.Shabestari Laser absorption spectrum 05-17-09 (23:55)
  • 274303 05/17/09 23:48 M.Shabestari Target---> strip charts
  • 274294 05/17/09 22:56 A. Camsonne Compton electron detector currents
  • 274293 05/17/09 22:52 A. Camsonne Lowered electron detector threshold to 11 during run 16954
  • 274280 05/17/09 21:22 LeRose HALL A TOOLS
  • 274279 05/17/09 21:19 LeRose Spot++
  • 274274 05/17/09 20:52 M.Shabestari Target--->strip charts
  • 274271 05/17/09 20:37 M.Shabestari Laser absorption spectrum 05-17-09 (20:35)
  • 274270 05/17/09 20:33 M.Shabestari NMR Frequency Sweep
  • 274254 05/17/09 19:58 hyde compton
  • 274249 05/17/09 19:14 miham ROC 5 problem during run #2376
  • 274237 05/17/09 17:56 M.Shabestari Target---> strip charts
  • 274236 05/17/09 17:52 Shabestari Laser absorption spectrum 05/1709 1750
  • 274235 05/17/09 17:33 LeRose Compton run 16952 started
  • 274230 05/17/09 17:01 adev BigBite WireChamber Threshold change!
  • 274227 05/17/09 16:58 miham New Thresholds for BB MWDC (5.5V) - Back To Production
  • 274222 05/17/09 16:51 miham New thresholds for BB MWDC (7.0V)
  • 274221 05/17/09 16:50 adev BigBite WireChamber Threshold change!
  • 274216 05/17/09 16:44 miham New thresholds for BB MWDC (6.5V)
  • 274211 05/17/09 16:38 miham New thresholds for BB MWDC (6.0V)
  • 274203 05/17/09 16:01 pking Day shift summary
  • 274202 05/17/09 15:42 Anderson NMR Measurement 62.1%
  • 274199 05/17/09 15:39 pking Started new HAPPEX run, 31387
  • 274192 05/17/09 14:48 pking Compton runs 16949 & 16950 had problems starting; 16951 is going.
  • 274190 05/17/09 14:41 pking Compton ROC5 crashed.
  • 274179 05/17/09 13:50 pking Started new Compton run, 16948, at 13:49
  • 274174 05/17/09 13:33 riordan BigBite Raw Trigger Asymmetries Near Spot Move
  • 274173 05/17/09 13:13 pking Hall A Tools screen
  • 274172 05/17/09 13:08 pking spot++ for run 2362
  • 274167 05/17/09 12:51 riordan Updated Golden ROOT files
  • 274166 05/17/09 12:33 R. Michaels same for R-HRS
  • 274161 05/17/09 12:28 Anderson Laser absorption spectrum 05/1709 1220 (Try again)
  • 274160 05/17/09 12:27 R. Michaels retune of parameter in pingpong script
  • 274159 05/17/09 12:14 Anderson Target Strip Charts 05_17_1215
  • 274158 05/17/09 12:11 Anderson Laser absorption spectrum 05/1709 1210
  • 274148 05/17/09 12:08 Anderson NMR Measurement 62.8%
  • 274143 05/17/09 11:50 pking Started new HAPPEX run, 31386
  • 274142 05/17/09 11:46 gjin gas shed: argon/ethane one bottle empty
  • 274141 05/17/09 11:44 pking Started new Compton run 16947 at 11:43.
  • 274124 05/17/09 10:02 M. Meziane ADC / TDC problem on the neutron detector: Left raw ACD/TDC 21
  • 274123 05/17/09 09:58 pking L-HRS DAQ stopped taking events in run 2354; two data disks are full!
  • 274122 05/17/09 09:48 M Meziane TDC Problem: Left Row TDC 21
  • 274116 05/17/09 09:34 pking Started Compton run 16946 at 09:34.
  • 274100 05/17/09 08:04 V. Sulkosky Owl Shift Summary, May 17, 2009
  • 274097 05/17/09 07:46 D. Parno NMR Measurement: 63.0%
  • 274093 05/17/09 07:18 D. Parno Compton restart (16945) and HAPPEX restart (31385)
  • 274088 05/17/09 07:10 V. Sulkosky Left HRS Stopped Counting for 2347 not 2447
  • 274087 05/17/09 06:52 V. Sulkosky Left HRS Stopped Counting for 2447
  • 274078 05/17/09 06:16 V. Sulkosky Scaler Rates at 10 uA and Deadtime
  • 274077 05/17/09 06:02 V. Sulkosky L-HRS Momentum to 2.277 GeV/c
  • 274072 05/17/09 05:38 V. Sulkosky Restarted the HAPPEX DAQ at 3:14 a.m. (New Run 31384)
  • 274071 05/17/09 05:33 V. Sulkosky Charge Asymmetry Chart
  • 274070 05/17/09 05:14 V. Sulkosky Rates During Raster Scan (late entry)
  • 274065 05/17/09 04:48 V. Sulkosky Final Harp Scan Before Production
  • 274060 05/17/09 04:29 W. Luo ND N1 TDC&ADC changed comparing to golden run
  • 274059 05/17/09 04:29 D. Parno Compton vertical scan
  • 274058 05/17/09 04:14 V. Sulkosky HAND HV's Off (21246)
  • 274044 05/17/09 04:02 V. Sulkosky Raster Check (Production)
  • 274043 05/17/09 04:00 D. Parno Laser absorption spectrum
  • 274041 05/17/09 03:59 D. Parno NMR Measurement: 63.7%
  • 274039 05/17/09 03:54 V. Sulkosky 6 x 6 Raster Check
  • 274034 05/17/09 03:50 V. Sulkosky 2 x 2 Raster Check
  • 274033 05/17/09 03:49 W. Luo run 2338: target BeO
  • 274030 05/17/09 03:43 D. Parno Target to polarized 3He
  • 274029 05/17/09 03:43 V. Sulkosky Spot check on BeO Target
  • 274026 05/17/09 03:35 D. Parno Target to Optics
  • 274023 05/17/09 03:30 V. Sulkosky Initial Harp Scans
  • 274022 05/17/09 03:24 V. Sulkosky Beam Back
  • 274021 05/17/09 03:20 adev BigBite WireChamber Measured Voltage change!
  • 274020 05/17/09 03:10 adev BigBite WireChamber Measured Voltage change!
  • 274019 05/17/09 02:57 V. Sulkosky Move Target to Empty
  • 274014 05/17/09 02:16 V. Sulkosky Still No Beam
  • 274013 05/17/09 01:33 Miha & Doug Summary of Threshold Change
  • 274012 05/17/09 01:25 Miha & Doug Summary of Tracking Changes
  • 274011 05/17/09 00:36 adev BigBite WireChamber Threshold change!
  • 274010 05/17/09 00:32 Miha & Doug No Beam So BigBite Back To 5.5V
  • 274009 05/17/09 00:12 Jin Huang Swing Shift Summary
  • 274007 05/17/09 00:07 Jin Huang Target: Rebooting Labview
  • Saturday

  • 274006 05/16/09 23:38 Miha & Doug BigBite Threshold Set To 6.5V
  • 274003 05/16/09 23:34 adev BigBite WireChamber Threshold change!
  • 273994 05/16/09 23:12 Jin Huang Hall A Tools
  • 273988 05/16/09 22:33 adev BigBite WireChamber Threshold change!
  • 273987 05/16/09 22:33 Miha & Doug BigBite Threshold Set To 5.5V
  • 273983 05/16/09 22:18 Kai Pan Gain of L-arm Cerenkov PMT6 dropped and recovered
  • 273982 05/16/09 22:12 riordan LHRS T2 Helicity Gated Scalars Double Pulsing?
  • 273979 05/16/09 21:45 riordan Set RHRS Angles in EPICS GUI
  • 273976 05/16/09 21:17 Jin Huang Spot Check
  • 273974 05/16/09 21:16 dwang NMR May 16, 21:13
  • 273960 05/16/09 20:54 Jin Huang Harp Scan Result
  • 273959 05/16/09 20:43 Jin Huang Beam Back
  • 273949 05/16/09 19:21 Jack (Techs) Right Q1
  • 273948 05/16/09 19:15 Jin Huang MCC called there may be another 1hour without beem
  • 273945 05/16/09 18:24 A. Camsonne Compton cavity sort of fixed
  • 273944 05/16/09 18:15 A. Camsonne Compton electron detector motion
  • 273943 05/16/09 17:55 Jin Huang RHRS Q1 is Stable
  • 273938 05/16/09 17:50 Jin Huang RHRS Optics Problem?
  • 273931 05/16/09 16:44 riordan RHRS Magnets Set
  • 273927 05/16/09 16:42 Jin Huang Scalar Screen shot after Q1 and ROC10 recovers
  • 273921 05/16/09 16:34 A. Camsonne Check with pulsed beam
  • 273919 05/16/09 16:20 Lindgren Shift summary
  • 273918 05/16/09 16:18 riordan RHRS Q1 Current RB Still 0
  • 273917 05/16/09 16:09 dwang NMR May 16, 16:06
  • 273916 05/16/09 16:00 hyde compton
  • 273914 05/16/09 15:49 puckett Yes, I do need to worry about "200% deadtime"--ROC10 is down
  • 273912 05/16/09 14:46 puckett "Tracks and Deadtime" page from onlineGUI gives bad result
  • 273911 05/16/09 14:26 lindgren spot++
  • 273906 05/16/09 14:19 Lindgren R-HRS Q1 readback current off
  • 273905 05/16/09 14:12 puckett online replay for run 2321/21228 looks fine despite right Q1 "zero current" readback
  • 273904 05/16/09 14:05 Lindgren Right Q1 problem
  • 273899 05/16/09 13:37 Anderson Target Strip Charts
  • 273898 05/16/09 13:33 lindgren Lecroy High voltage
  • 273897 05/16/09 13:32 Anderson Laser absorption spectrum
  • 273896 05/16/09 13:30 Lindgren charge asymmetry current strip and strip charts
  • 273895 05/16/09 13:28 lindgren hall A tools
  • 273886 05/16/09 12:15 Anderson NMR Measurement 62.2%
  • 273872 05/16/09 11:37 R. Lindgren Happex run ended
  • 273867 05/16/09 11:15 R. Lindgren production running update
  • 273854 05/16/09 09:28 puckett online replay for 2313/21220 after restarting CODA looks normal
  • 273853 05/16/09 09:19 R. Lindgren coda crash
  • 273852 05/16/09 09:13 A. Puckett CODA hangs during run 2312/21219, kcoda/startcoda fixed things
  • 273845 05/16/09 08:36 R. Lindgren Day shift beginning
  • 273839 05/16/09 08:05 Ibrahim Owl Shift Summary (May 16, 2009)
  • 273838 05/16/09 08:02 Anderson NMR May 16, 0800
  • 273825 05/16/09 06:27 Ibrahim Restarted the HAPPEX DAQ at 6:23 a.m. (New Run 31381)
  • 273812 05/16/09 05:07 V. Sulkosky VDC Wire Plots
  • 273811 05/16/09 05:06 D. Parno Laser absorption spectrum
  • 273810 05/16/09 05:06 V. Sulkosky VDC Efficiency Plots
  • 273809 05/16/09 04:55 V. Sulkosky Run 2302 Stopped Counting
  • 273802 05/16/09 04:12 D. Parno NMR Measurement: 61.7%
  • 273790 05/16/09 04:01 Ibrahim Spot++ for the Checklist (Run 2302)
  • 273781 05/16/09 02:41 Ibrahim Restarted the HAPPEX DAQ (New Run 31380)
  • 273780 05/16/09 02:41 V. Sulkosky Trigger Rates at 10 uA
  • 273767 05/16/09 01:17 D. Parno Compton Polarization History May 9 - 15
  • 273762 05/16/09 00:29 D. Parno NMR Measurement: 61.7%
  • 273757 05/16/09 00:05 miham Lumi replay plots for runs #2282 - #2293
  • 273756 05/16/09 00:03 pking Swing shift summary
  • Friday

  • 273742 05/15/09 22:27 miham Strange LUMI replay plots
  • 273741 05/15/09 22:27 pking Started new HAPPEX run, 31379
  • 273736 05/15/09 22:00 miham New golden run for Right arm and ND #21198
  • 273731 05/15/09 21:21 A. Camsonne Cannot lock Compton laser cavity
  • 273730 05/15/09 21:20 A. Camsonne Cannot lock Compton laser cavity
  • 273723 05/15/09 20:27 dwang NMR May 15, 20:25
  • 273711 05/15/09 20:05 pking Run 21194 had R-HRS momentum set to 2.225
  • 273706 05/15/09 19:45 pking Tools screen for checklist
  • 273705 05/15/09 19:32 pking spot++ in run 2287: now with screenshots
  • 273704 05/15/09 19:31 pking spot++ in run 2287
  • 273703 05/15/09 19:18 pking HA Tools after R-HRS momentum change
  • 273700 05/15/09 19:08 Moffit New golden run for bigbite replay plots..
  • 273699 05/15/09 19:06 pking Changing the R-HRS momentum
  • 273696 05/15/09 18:34 Moffit db_BB.tp.dat modifed..
  • 273691 05/15/09 18:19 Moffit RHRS EDTM on
  • 273690 05/15/09 17:52 pking Started new HAPPEX run, 31378
  • 273681 05/15/09 16:54 M.Meziane Deadtime issue?? disappeared
  • 273676 05/15/09 16:33 M.Meziane RUN# 21189: Shower Total Energy Divided by Momemntum histogram does not match the golden run
  • 273675 05/15/09 16:25 M Meziane Dead time issue??
  • 273670 05/15/09 16:08 dwang NMR May 15, 16:00
  • 273668 05/15/09 16:05 gjin day shift summary
  • 273662 05/15/09 15:15 Katich target temp swing
  • 273653 05/15/09 14:24 Yawei It is OK.
  • 273652 05/15/09 14:23 gjin hallA tools screen
  • 273651 05/15/09 14:22 gjin BCM asymmetry in Happex
  • 273650 05/15/09 14:22 ellie Errorrs in do_all and show_all for RHRS
  • 273649 05/15/09 14:21 gjin charge asymmetry too wide
  • 273644 05/15/09 13:56 gjin replay_R does not work
  • 273635 05/15/09 13:00 gjin spot++
  • 273621 05/15/09 12:05 Anderson NMR Run 5/15/09 1200
  • 273619 05/15/09 12:01 riordan Spot Change
  • 273607 05/15/09 10:53 Moffit Re: Replay on 2270
  • 273606 05/15/09 10:42 Norum Replay 2270
  • 273601 05/15/09 10:26 Anderson Target Temperature Swing
  • 273595 05/15/09 10:11 Norum Run 2266
  • 273594 05/15/09 10:06 Norum BB MWDC hits/wire
  • 273579 05/15/09 09:06 Norum ROC sync check
  • 273576 05/15/09 08:56 gjin compton: is it normal? run 16915
  • 273568 05/15/09 08:09 V. Sulkosky Owl Shift Summary, May 15, 2009
  • 273567 05/15/09 08:04 V. Sulkosky Offline Deadtime for 2252 and 21159
  • 273565 05/15/09 07:52 Holmstrom NMR Measurement: 62.0%
  • 273561 05/15/09 07:30 V. Sulkosky Hall A Tools
  • 273556 05/15/09 06:59 D. Parno Compton Cavity Having Some Trouble
  • 273555 05/15/09 06:51 D. Parno HAPPEX Restart (31375); Compton restart (16915)
  • 273546 05/15/09 06:08 D. Parno Anomaly in LHRS Cerenkov ADC 6
  • 273541 05/15/09 05:39 D. Parno VDC Efficiency Captures
  • 273540 05/15/09 05:36 V. Sulkosky Updated Hall C Current Variable for Epics
  • 273535 05/15/09 05:02 D. Parno Compton restart (16914)
  • 273526 05/15/09 04:08 V. Sulkosky Re: Updated Offline Deadtime Macro (LHRS)
  • 273516 05/15/09 04:05 Updated Offline Deadtime Macro (LHRS)
  • 273515 05/15/09 03:49 Holmstrom NMR Measurement: 62.6%
  • 273510 05/15/09 03:24 V. Sulkosky Added Physics Replay to R-HRS
  • 273505 05/15/09 02:49 D. Parno HAPPEX restart (31374), Compton restart (16913)
  • 273500 05/15/09 02:36 V. Sulkosky Set LHRS/BB Golden Run to 2250, Set RHRS/HAND Golden Run to 21157
  • 273495 05/15/09 02:01 D. Parno RHRS Online Replay: Total Shower E/P graph missing
  • 273494 05/15/09 01:53 V. Sulkosky Lumi's After HV Adjustment
  • 273493 05/15/09 01:44 V. Sulkosky New Lumi HV's
  • 273488 05/15/09 01:28 V. Sulkosky Lumi's 1, 6 and 7 Saturated at 10 uA
  • 273487 05/15/09 01:09 D. Parno Compton DAQ restarted for Run 16912
  • 273486 05/15/09 01:07 D. Parno Rates at 10 uA
  • 273485 05/15/09 01:06 V. Sulkosky Beam at 10 uA
  • 273480 05/15/09 00:33 ellie Set LHRS/BB Golden Run to 2248, Set RHRS/HAND Golden Run to 21155
  • 273475 05/15/09 00:01 pking Swing shift summary
  • Thursday

  • 273471 05/14/09 23:48 ellie Changed BB HV from -1550 to -1525
  • 273470 05/14/09 23:47 Zhihong Ye 2rd NMR_05_14_swing_shift, Pol=63.208%
  • 273467 05/14/09 23:41 ellie dE TDC Histograms Before and After Threshold Change
  • 273464 05/14/09 23:10 ellie Changed MWDC_Front Threshold to 4.5V (was 4.0V)
  • 273461 05/14/09 22:53 ellie HAPPEX Restarted (31373), Compton Restarted (16911)
  • 273460 05/14/09 22:47 ellie VDC Efficiency Screen Grabs
  • 273459 05/14/09 22:45 ellie BPM looks better on new run
  • 273458 05/14/09 22:41 riordan spot++ 2245 looks weird
  • 273457 05/14/09 22:35 ellie Strange BPM. Beam trip? Looks weird.
  • 273450 05/14/09 22:17 ellie BB dE Threshold set to 100 (was 150)
  • 273447 05/14/09 22:06 Zhihong Ye Target Graph 1
  • 273446 05/14/09 22:04 Zhihong Ye Target Graph 1
  • 273445 05/14/09 21:46 pking CODA stopped being able to write events at end of 2243; data disk 4 was full.
  • 273440 05/14/09 21:23 ellie Lumi vs. BCM momentarily weird
  • 273439 05/14/09 21:14 M. Meziane ps3=3 LHRS
  • 273438 05/14/09 21:13 pking Late entry: Harp measurement after Moller backout
  • 273437 05/14/09 21:09 ellie RHRS Preshower ADC 13 returns!
  • 273432 05/14/09 20:59 ellie BB Bars Hit, Threshold = 400
  • 273431 05/14/09 20:49 ellie Scalers Screen Grab
  • 273430 05/14/09 20:47 ellie Compton Rebooted, Started Run 16910
  • 273429 05/14/09 20:39 ellie Compton End of Run Failed, Rebooting CODA
  • 273428 05/14/09 20:37 ellie HAPPEX Restarted, Current Run: 31372
  • 273425 05/14/09 20:35 ellie Changed BB E Threshold to 300
  • 273422 05/14/09 20:27 ellie Scalars Screen Grab
  • 273421 05/14/09 20:23 pking Helium target cell was in beam beginning for runs (2235, 21142).
  • 273420 05/14/09 20:10 pking Tools screen
  • 273407 05/14/09 20:01 pking Runs (2240, 21147) should be marked as production on He-3 target
  • 273406 05/14/09 19:49 pking Spot++ for run 2240
  • 273401 05/14/09 19:35 pking Okay, the hatsv3:2005 HV was off and stuck; someone pushed HV OFF?
  • 273394 05/14/09 19:15 Bo Re:R-HRS trigger rates are zero
  • 273393 05/14/09 19:12 pking R-HRS rates still zero after coda restart.
  • 273386 05/14/09 19:01 pking R-HRS trigger rates are zero
  • 273379 05/14/09 18:50 Zhihong Ye NMR_05_14_swing_shift, Pol=63.5931%
  • 273378 05/14/09 18:47 pking spott from run 2234
  • 273375 05/14/09 18:43 pking spot++ in run 2233
  • 273368 05/14/09 18:32 pking Optics target in; request 2 uA raster off.
  • 273367 05/14/09 18:32 Zhihong Ye Target moves to Optics
  • 273366 05/14/09 18:27 A. Saha Moller Measurement
  • 273365 05/14/09 17:53 riordan BB dE/E Plane Thresholds Restored
  • 273361 05/14/09 16:03 LeRose Shift Summary (Day)
  • 273354 05/14/09 15:56 Higinbotham Test Runs Before Moller Quads Off
  • 273353 05/14/09 15:23 LeRose Beginning Moller Measurement
  • 273349 05/14/09 15:14 miham Changed back the BB threshold voltages
  • 273346 05/14/09 15:11 B. Norum 2nd_NMR_05_14_day_shift
  • 273340 05/14/09 14:56 LeRose Spot++
  • 273337 05/14/09 14:37 Jack (Techs) The logbook does not like the character ASCII 60 (decimal)
  • 273336 05/14/09 14:31 Bo wide pedestal preshower channels (5R and 18R) in RHRS
  • 273335 05/14/09 14:30 Jack (Techs) Right Q1 interlocks modified
  • 273332 05/14/09 14:24 LeRose Beam Permit
  • 273331 05/14/09 14:23 Jack (Techs) Right and Left Q1s polarity readback changed
  • 273330 05/14/09 14:19 Bo wide pedestal in preshower (5R and 18R)
  • 273329 05/14/09 14:13 LeRose Power Permit
  • 273320 05/14/09 14:02 Moffit ROC8 missing gate problem... solved?
  • 273316 05/14/09 13:28 miham New Threshold values for the BB trigger planes
  • 273313 05/14/09 13:20 V. Sulkosky Re: Corrected the T2 and T5 in the HRS-L
  • 273312 05/14/09 13:09 miham Corrected the T2 and T5 in the HRS-L
  • 273311 05/14/09 13:08 LeRose Beam Returning
  • 273308 05/14/09 12:49 Yi Qiang New settings for Q-wave plates
  • 273292 05/14/09 12:03 R. Subedi access activity (PVDIS work)
  • 273291 05/14/09 12:03 Yi Qiang Target Strip Charts of the day
  • 273287 05/14/09 11:49 Ramesh/Bo Preshower Block 14 signal cable bad, replaced
  • 273286 05/14/09 11:42 R. Michaels power-cycled Compton FADC crate, PORT6 of hareboot4
  • 273281 05/14/09 11:37 Yi Qiang NMR Measurment: 62.7%
  • 273274 05/14/09 10:22 gjin hit efficiency of run 2180
  • 273273 05/14/09 10:10 gjin track efficiency results for run 2207-2211
  • 273272 05/14/09 10:06 Moffit Re: updated timerocs
  • 273271 05/14/09 09:47 Moffit updated timerocs for RightHRS
  • 273270 05/14/09 08:50 LeRose Controlled Access
  • 273269 05/14/09 08:45 Yi Qiang Only two RTDs on Reference cell are working after the cell change
  • 273268 05/14/09 08:29 LeRose Controlled Access
  • 273266 05/14/09 08:00 Holmstrom Owl Shift Summary
  • 273263 05/14/09 07:23 miham Ref. Cell RTD problem
  • 273262 05/14/09 07:13 miham No space left on the device!!!
  • 273253 05/14/09 06:22 miham BB Wire Hit plots
  • 273250 05/14/09 05:47 miham Pumping laser Absorption Spectrum
  • 273249 05/14/09 05:38 miham NMR Frequency sweep - Pol.: 61.96%
  • 273242 05/14/09 04:32 miham New threshold for the dE-TDC
  • 273228 05/14/09 03:53 miham MWDC wire with LOW number of hits
  • 273223 05/14/09 03:22 miham BB MWDC HV scan
  • 273214 05/14/09 02:31 miham Spot++ check during run #2212
  • 273211 05/14/09 02:18 miham BB MWDC HV changed BACK to 1550V
  • 273206 05/14/09 02:07 miham BB MWDC HV changed to 1350V
  • 273203 05/14/09 02:03 miham Right arm DAQ problem
  • 273197 05/14/09 01:50 miham BB MWDC HV changed to 1400V
  • 273192 05/14/09 01:40 miham BB MWDC HV changed to 1450V
  • 273187 05/14/09 01:30 miham BB MWDC HV changed to 1500V
  • 273180 05/14/09 00:18 T. Holmstrom NMR: 61.9%
  • Wednesday

  • 273174 05/13/09 23:56 pking Swing shift summary
  • 273169 05/13/09 23:28 ellie HAPPEX Restarted, Current Run: 31368
  • 273164 05/13/09 22:52 adaq LHRS Cerenkov ADC 6 Low, Beam Trip?
  • 273151 05/13/09 21:19 ellie VDC Efficiency Screen Grabs
  • 273148 05/13/09 21:09 ellie LHRS Momentum Transfer, Energy Transfer, and Nu vs x are back!
  • 273145 05/13/09 21:08 pking Late entry: Compton computer is frozen
  • 273143 05/13/09 20:40 A. Deur NMR: 61.6%
  • 273134 05/13/09 20:08 ellie Set LHRS/BB Golden Run to 2194
  • 273124 05/13/09 19:56 ellie BB Replay: HALLA_p{abs(HALLA_p-1200)<500} Histogram Empty
  • 273123 05/13/09 19:55 pking Hall A tools screen
  • 273122 05/13/09 19:54 pking spot++ raster size check
  • 273121 05/13/09 19:53 ellie LHRS missing Mom. Transfer, Energy Transfer, and Nu vs. x
  • 273120 05/13/09 19:49 V. Sulkosky Re: RHRS Shower ADCs off?
  • 273115 05/13/09 19:25 ellie Set RHRS/HAND Golden Run to 21094
  • 273114 05/13/09 19:24 gjin MWDC hit efficiency of run 2170
  • 273113 05/13/09 19:22 ellie RHRS Shower ADCs off?
  • 273112 05/13/09 19:21 gjin hit efficiency of run 2154
  • 273107 05/13/09 18:44 pking R-HRS has been ramped up during runs (2193, 21093)
  • 273102 05/13/09 18:22 pking R-HRS Q1 won't ramp up.
  • 273099 05/13/09 18:08 pking R-HRS Q1 has ramped down.
  • 273088 05/13/09 17:11 brads BCM calibration coefficients for "Tools screen" restored
  • 273087 05/13/09 17:02 Jin Huang Spin signals Test
  • 273078 05/13/09 16:37 A. Deur NMR: 62.6%
  • 273075 05/13/09 16:35 pking Started new Compton run 16894 at 16:30.
  • 273072 05/13/09 16:25 A. Deur Screen shot of target temperatures, B-field and Pt.
  • 273071 05/13/09 16:23 A. Deur Laser absorption spectrum
  • 273069 05/13/09 16:04 averett day shift summary
  • 273060 05/13/09 15:24 mcanan target_screen_shots
  • 273058 05/13/09 15:02 adaq Hall A tools screen shot
  • 273054 05/13/09 14:27 Yi Qiang Late Entry: EPR calibrations of cell MOSS
  • 273053 05/13/09 14:25 Higinbotham How I learned to stop worring and love the coil.
  • 273051 05/13/09 14:21 deng Lumi HV adjusted for beam 7.5uA--continue
  • 273049 05/13/09 14:20 deng Lumi HV adjusted for beam 7.5 uA
  • 273047 05/13/09 14:17 R. Michaels updated datamon for L-HRS
  • 273042 05/13/09 14:10 gjin changed db_BB.mwdc (the time window) & replay_L(add optics)
  • 273041 05/13/09 14:08 Yi Qiang Late entry: current settings for rotatable Q-wave plates of longitudianl laser
  • 273032 05/13/09 13:17 averett lumi 7 problem
  • 273027 05/13/09 12:40 mcanan attachment
  • 273022 05/13/09 12:37 mcanan laser_absorption_spectrum
  • 273021 05/13/09 12:26 mcanan 2nd_NMR_05_13_day_shift
  • 273016 05/13/09 12:08 A. Camsonne Compton cavity did not lock
  • 273002 05/13/09 11:47 gjin MWDC HV and threshold scan on May 12th swing shift
  • 272997 05/13/09 11:26 V. Sulkosky Re: Updated End-of-Run Deadtime Calculation L-HRS
  • 272980 05/13/09 10:38 R. Michaels re: datamon
  • 272979 05/13/09 10:26 V. Sulkosky Re: Right arm gas Cherenkov, Ch2 gain low?
  • 272978 05/13/09 10:13 Higinbotham Left HRS Datamon Miscalculating?
  • 272973 05/13/09 09:06 A. Tobias Happex start new run 31365
  • 272970 05/13/09 09:05 miham New HVsetting for MWDC
  • 272969 05/13/09 09:04 A. Tobias Synchronize running L-arm and R-arm
  • 272968 05/13/09 09:00 R. Michaels Compton FADC crate dead
  • 272965 05/13/09 08:58 averett spot ++
  • 272962 05/13/09 08:52 mcanan NMR_05_13_day_shift
  • 272961 05/13/09 08:48 mcanan Target Motion
  • 272956 05/13/09 08:32 A. Tobias Owl Shift Summary
  • 272946 05/13/09 07:07 A. Tobias Laser absortion spectrum OWL shift
  • 272945 05/13/09 06:48 A. Tobias Hall A Tools snapshot
  • 272944 05/13/09 06:48 miham Adaq computers 4-8 extremely slow
  • 272943 05/13/09 06:40 miham BB & HRS-L DAQ test in unbuffered mode
  • 272941 05/13/09 06:32 A. Tobias spot++ D2 reference cell 4x4mm
  • 272939 05/13/09 06:29 miham HAPPEX daq problems
  • 272938 05/13/09 05:52 A. Tobias HAPPEX ROC26 not responding - reboot & start new HAPPEX run#31364 at 5:25
  • 272932 05/13/09 04:54 miham New Threshold values for BB
  • 272926 05/13/09 04:30 A. Tobias HACSBC2 Keyboard Frozen
  • 272925 05/13/09 04:27 A. Tobias Rollup Door is Open down in Hall
  • 272915 05/13/09 03:55 miham 7th HV change for BB Scintil. planes
  • 272910 05/13/09 03:05 miham 6th HV change for BB Scintil. planes
  • 272907 05/13/09 02:56 miham New Thresholds for BB
  • 272904 05/13/09 02:42 A. Tobias Compton continued ...
  • 272903 05/13/09 02:35 A. Tobias Compton not running properly
  • 272901 05/13/09 02:25 jpchen late runs are fine
  • 272899 05/13/09 01:59 A. Tobias Compton Diode Laser was not ON
  • 272898 05/13/09 01:44 miham 5th HV change for BB Scintil. planes
  • 272897 05/13/09 01:36 miham dE/E plots on deuterium target.
  • 272894 05/13/09 01:15 miham Elastic deuterons in HRS-L
  • 272893 05/13/09 01:04 miham dE/E plots after 4rd iteration of BB gain maching (Made on Sunday)
  • 272890 05/13/09 00:41 ellie Still More HAND HV Tweaking
  • 272887 05/13/09 00:30 A. Tobias Restart HAPPEX again: (31361 start)
  • 272886 05/13/09 00:28 jpchen swing shift summary
  • 272883 05/13/09 00:22 A. Tobias Stop and Restart HAPPEX run (31359 stop & 31360 start)
  • Tuesday

  • 272879 05/12/09 23:52 ellie More HAND HV Tweaks
  • 272878 05/12/09 23:52 ellie More HAND HV Tweaks
  • 272874 05/12/09 23:36 jpchen BB hits per wire
  • 272869 05/12/09 23:30 jpchen Left arm VDC hit: many zeros
  • 272868 05/12/09 23:28 jpchen left arm with multi 12C+BeO
  • 272864 05/12/09 23:21 jpchen Multi-12C, Left arm: OK
  • 272863 05/12/09 23:11 V. Sulkosky Re: L-HRS T6 [--] Counting More
  • 272862 05/12/09 23:04 gjin run2141-2149: MWDC scan
  • 272861 05/12/09 22:53 jpchen Right arm gas Cherenkov, Ch2 gain low?
  • 272860 05/12/09 22:51 jpchen right preshower Ch13 no signal
  • 272859 05/12/09 22:51 ellie Some Small HAND HV Tweaks
  • 272854 05/12/09 22:45 jpchen right arm y_target, +side window blocked
  • 272850 05/12/09 22:40 V. Sulkosky L-HRS T6 [--] Counting More
  • 272836 05/12/09 22:01 jpchen BB check-out with D2 target now
  • 272835 05/12/09 21:55 jpchen Data taking for multi-12C
  • 272834 05/12/09 21:47 jpchen center beam on reference cell and pol. He3 cell
  • 272828 05/12/09 21:19 Yawei filled D2 into reference cell to 100psig
  • 272826 05/12/09 21:12 Yawei Target moves to reference cell:D2
  • 272813 05/12/09 20:40 Yawei Trigger rates
  • 272811 05/12/09 20:32 V. Sulkosky Updated End-of-Run Deadtime Calculation L-HRS
  • 272810 05/12/09 20:31 V. Sulkosky Updated End-of-Run Deadtime Calculation L-HRS
  • 272809 05/12/09 20:31 Yawei Target moves to BeO
  • 272808 05/12/09 20:30 Yawei spot++ check, 3He, 6x6, x=-0.5
  • 272803 05/12/09 20:21 Yawei spot++ check, 6x6,3He target, x=-1
  • 272787 05/12/09 20:13 Yawei spot++ check, 3He, 4x4
  • 272784 05/12/09 20:07 Yawei Target moves to Polarized 3He
  • 272781 05/12/09 20:02 V. Sulkosky Re: T2 and T5 Swapped in L-HRS
  • 272780 05/12/09 20:02 Yawei spot++, 4x4, x=-1
  • 272777 05/12/09 19:56 Yawei spot++ check,6x6, x=-1
  • 272772 05/12/09 19:52 Yawei spot++ check, 6x6, x=0
  • 272769 05/12/09 19:49 V. Sulkosky Checked Afterpulsing from Multi-hit TDC
  • 272766 05/12/09 19:46 Yawei spot++ check,6x6
  • 272763 05/12/09 19:38 V. Sulkosky Still See Missing Synch events for ROC 8
  • 272758 05/12/09 19:27 jpchen reference cell (D2): raster 4x4--> 6x6 spot++
  • 272751 05/12/09 19:17 riordan Neutron Arm and BB DC On
  • 272750 05/12/09 19:13 jpchen reference cell filled with 1 atm D2 gas, moved target to reference cell
  • 272749 05/12/09 19:12 Yawei Target moves to Ref Cell:D2
  • 272746 05/12/09 19:09 jpchen reference cell: pressure reading 3 psig off, reset
  • 272745 05/12/09 19:05 jpchen raster check: 4x4 (MCC 7x5): good
  • 272743 05/12/09 19:00 jpchen hapr scan: good
  • 272738 05/12/09 18:57 jpchen spot check with raster off
  • 272735 05/12/09 18:54 V. Sulkosky/ R-HRS T6 Trigger From BigBite
  • 272733 05/12/09 18:52 riordan Updated Alarm Handler to Include RHRS
  • 272730 05/12/09 18:49 V. Sulkosky ND and MWDC Off for Beam Tuning
  • 272729 05/12/09 18:49 jpchen Hall activities completed, back to beam permit. beam back
  • 272728 05/12/09 18:48 Yawei Target moves to Optics
  • 272727 05/12/09 18:44 V. Sulkosky T2 and T5 Swapped in L-HRS
  • 272726 05/12/09 18:19 miham New BB threshold Values
  • 272725 05/12/09 18:18 riordan RHRS Top VDC HV in Start of Run
  • 272724 05/12/09 18:14 V. Sulkosky R-HRS S1 and S2 TDC Spectra
  • 272723 05/12/09 17:59 V. Sulkosky R-HRS S1 and S2 PMTs Gain Matched
  • 272722 05/12/09 17:44 V. Sulkosky Disabled Golden Root Files for L-HRS and R-HRS Replays
  • 272719 05/12/09 17:26 Yi Qiang Field gradient test
  • 272713 05/12/09 16:29 A. Camsonne Electron detector motion GUI test done
  • 272708 05/12/09 15:42 hyde compton
  • 272706 05/12/09 15:13 techs Checklist
  • 272705 05/12/09 15:12 L. Zhu target: P=51.1%
  • 272704 05/12/09 14:54 A. Camsonne Electron detector safety features test
  • 272703 05/12/09 14:07 ellie Adjusted HAND HV
  • 272701 05/12/09 13:51 techs hall A beamline valves
  • 272692 05/12/09 12:29 riordan Left and Right HRS Set
  • 272682 05/12/09 12:03 L. Zhu target: P=45.8%
  • 272680 05/12/09 11:59 V. Sulkosky Moved R-HRS VDC TOP HV
  • 272679 05/12/09 11:56 Yi Qiang Connected Deuteron bottle to the H2 inlet for the reference
  • 272677 05/12/09 10:55 L. Zhu target: P=37.8%
  • 272673 05/12/09 09:58 L. Zhu R-arm S1/S2 ADC
  • 272672 05/12/09 09:48 Ramesh/Xiaoc access activity (PVDIS work)
  • 272670 05/12/09 09:47 bogdanw HRS-R S1/S2 summary
  • 272668 05/12/09 09:09 Moffit BB TS debug port plugged in
  • 272667 05/12/09 09:08 Moffit Added Trigger Cable termination to ROC8 TI
  • 272665 05/12/09 08:03 miham OWL Shift summary (05/12/09)
  • 272663 05/12/09 07:20 Fansler/Ewin Spectometers/Big Bite
  • 272662 05/12/09 06:00 Target Group Target Activities
  • 272661 05/12/09 05:57 Jin/Yi NMR spin up started
  • 272660 05/12/09 05:34 Yi/Jin Longitudinal Quater Wave Plate Calibration
  • 272648 05/12/09 03:44 Yi Qiang Laser and fiber mapping for Longitudianl pumping
  • 272647 05/12/09 03:38 Yi Qiang Fiber mappings for Longitudinal pumping
  • Monday

  • 272644 05/11/09 22:44 ellie Updated HAND HV
  • 272643 05/11/09 22:44 ellie Updated HAND HV
  • 272631 05/11/09 20:07 Ramesh/Xiaoc Access activity (PVDIS work)
  • 272630 05/11/09 19:43 Bo/Vince Update on checking ROC8 busy signal
  • 272629 05/11/09 19:37 ellie New HAND HV
  • 272628 05/11/09 19:36 V. Sulkosky Re: BB triggers to RHRS on long ribbon cable (grey band)
  • 272627 05/11/09 19:21 ellie HAND HV Updated, Cosmics Run Started
  • 272622 05/11/09 17:55 R. Michaels scaler moved to Compton FADC crate
  • 272621 05/11/09 17:52 ellie Beginning long HAND HV update
  • 272620 05/11/09 17:44 X. Jiang Two-Arm DAQ-sync was only required by the Ay experiment
  • 272619 05/11/09 17:44 X. Jiang Two-Arm DAQ-sync was only required by the Ay experiment
  • 272617 05/11/09 16:48 M.Shabestrai run 21025
  • 272615 05/11/09 16:41 bogdanw HRS S1/S2 counters update
  • 272614 05/11/09 16:41 M.Shabestrai Run 21025
  • 272610 05/11/09 16:04 miham Problem with Dead TDC channels in BB dE-plane hopefully resolved
  • 272609 05/11/09 15:53 hyde Compton Vacuum
  • 272608 05/11/09 15:45 V. Sulkosky Re: Instruction to synchronize running L-arm and R-arm
  • 272603 05/11/09 15:24 Bo Instruction to synchronize running L-arm and R-arm
  • 272602 05/11/09 15:20 ellie RHRS CODA on ROC8Only configuration, Getting Cosmics Data for HAND
  • 272599 05/11/09 15:18 V. Sulkosky Re: BB triggers to RHRS on long ribbon cable (grey band)
  • 272594 05/11/09 15:00 Yawei Grading for the Fiber Tips(new)
  • 272593 05/11/09 14:54 Yawei Grading for the Fiber Tips(new)
  • 272592 05/11/09 14:38 Yawei Grading for the Fiber Tips(new)
  • 272590 05/11/09 14:05 R. Michaels scaler in compton FADC crate
  • 272572 05/11/09 12:43 V. Sulkosky/ ROC 8 BUSY Working Now
  • 272568 05/11/09 12:34 gjin MWDC HV and threshold scan (results are preliminary)
  • 272555 05/11/09 12:12 Moffit BB triggers to RHRS on long ribbon cable (grey band)
  • 272554 05/11/09 11:33 Yi Qiang Target holding field is set to Longitudinal +
  • 272553 05/11/09 11:28 Moffit No golden rootfile for HAND analysis
  • 272551 05/11/09 11:06 R. Michaels re: R-HRS edtm
  • 272546 05/11/09 10:45 Moffit Missing TDC signals from N1-13L
  • 272545 05/11/09 10:41 Moffit Figures for new HAND timing config (p.1)
  • 272543 05/11/09 10:31 Moffit New HAND setup... p.1
  • 272541 05/11/09 10:07 Moffit HAND edtm signals, present configuration (same as was for Ay and e'n)
  • 272540 05/11/09 10:01 mcanan Target Motion
  • 272539 05/11/09 09:47 V. Sulkosky Re: RightHRS edtm
  • 272537 05/11/09 09:41 mcanan Target Motion
  • 272536 05/11/09 09:40 mcanan Reference Cell
  • 272535 05/11/09 09:36 Bo Run L-arm and R-arm independently now
  • 272533 05/11/09 09:28 atrig Trigger-Download
  • 272531 05/11/09 09:24 mcanan Target Motion
  • 272527 05/11/09 09:20 jpchen Day shift summary
  • 272526 05/11/09 09:13 jpchen planed hall activities
  • 272525 05/11/09 09:05 Moffit Re: RightHRS edtm
  • 272524 05/11/09 08:59 jpchen Hall will be in restricted access
  • 272523 05/11/09 08:56 jpchen Lasers off/locked out. oven off
  • 272522 05/11/09 08:54 jpchen End of production for E05-015/E08-005.
  • 272519 05/11/09 08:46 Moffit RightHRS edtm on.
  • 272514 05/11/09 08:03 A. Tobias Owl Shift Summary 11may2009
  • 272513 05/11/09 07:42 miham laser absorption spectrum
  • 272512 05/11/09 07:35 A. Tobias Right Arm VDCs tripped - running LHRS-only production on 3He
  • 272508 05/11/09 07:17 miham Big HV trip in ND.
  • 272507 05/11/09 07:07 miham Spot++ check for (6x6)mm raster on 3He
  • 272502 05/11/09 07:04 miham Spot++ check for (2x2mm) raster on 3He
  • 272499 05/11/09 06:56 miham Target moved to pol. 3He Target
  • 272497 05/11/09 06:53 miham spot++ 4x4 raster (7x5mcc) on Optics 2microA
  • 272494 05/11/09 06:47 miham Target moved to Optics Target
  • 272492 05/11/09 06:36 A. Tobias HARP Scans after beam recovery
  • 272491 05/11/09 06:11 miham Target moved to Empty Target
  • 272490 05/11/09 06:00 miham Target moved to Optics
  • 272489 05/11/09 05:56 Target target: Spin Flip Report: Target Halog report - Target Empty - Beam is comming back
  • 272488 05/11/09 05:24 A. Tobias Hall A Tools Screen Shot - No beam
  • 272478 05/11/09 00:17 Al Tobias Target Moved to Empty & "Auto Flip Set" is OFF
  • 272477 05/11/09 00:14 Yi Qiang Experiment Summary for May 10, 2009
  • 272475 05/11/09 00:10 R. Michaels compton files not splitting
  • 272474 05/11/09 00:02 R. Michaels target to empty
  • Sunday

  • 272473 05/10/09 23:56 Markowitz Swing shift summary
  • 272469 05/10/09 22:39 V. Sulkosky Re: ADCs for run 2096; are fine
  • 272468 05/10/09 22:05 Zhihong Ye target graphs
  • 272467 05/10/09 22:04 Zhiho Target Control Screen Shot
  • 272458 05/10/09 20:25 Mitra S. ADCs for run 2096; are they OK?
  • 272449 05/10/09 19:46 Markowitz raster screen
  • 272444 05/10/09 19:29 Chiranjib Right arm VDC efficiencies
  • 272443 05/10/09 19:28 Chiranjib Left arm VDC efficiencies
  • 272436 05/10/09 18:10 Zhiho Target moves to He3 Cell
  • 272433 05/10/09 17:31 adaq tools screenshot
  • 272428 05/10/09 17:11 Markowitz raster off
  • 272423 05/10/09 17:06 Chiranjib Carbon foils/ Be0 for the optics runs 2090(Left) and 20996(Right)
  • 272414 05/10/09 16:13 miham Next HV change for BB Scintil. planes
  • 272412 05/10/09 16:10 Lindgren Day Shift Summary
  • 272411 05/10/09 15:52 Lindgren Optics test
  • 272408 05/10/09 15:48 Lindgren Optics target
  • 272407 05/10/09 15:41 chench move the target from the 3He cell to optics
  • 272404 05/10/09 15:28 Target target: Spin Flip Report: spin flit system
  • 272403 05/10/09 15:25 chench screen shot laser absorption spectrum
  • 272402 05/10/09 15:13 W. Luo run 2089: L-arm Cerenkov ADC channel 6 gain drop
  • 272401 05/10/09 15:09 Yi Qiang Auto spin flip was turned ON before run 2089/20995
  • 272400 05/10/09 15:07 Lindgren BB Test Completed
  • 272399 05/10/09 15:07 chench Target Status
  • 272398 05/10/09 15:03 lindgren happex
  • 272397 05/10/09 15:01 lindgren Lecroy voltage
  • 272396 05/10/09 14:58 lindgren energy strip
  • 272393 05/10/09 14:52 chench Flip the spin of the target and turn off the auto flip
  • 272392 05/10/09 14:50 lindgren ozone strip and compton strip
  • 272389 05/10/09 14:45 lindgren charge asymmetry
  • 272388 05/10/09 14:43 lindgren tools
  • 272371 05/10/09 13:47 miham dE/E plots after 3rd iteration of BB gain maching
  • 272339 05/10/09 12:31 Lindgren BigBite
  • 272326 05/10/09 10:48 Lindgren DAQ Crash
  • 272322 05/10/09 10:42 Lindgren Back to Production Running
  • 272317 05/10/09 10:27 miham HV mapping table for the 3rd iteration of Gain Matching (this time printed correctly)
  • 272316 05/10/09 10:20 Lindgren Beam back
  • 272315 05/10/09 10:14 Lindgren CHL problem
  • 272314 05/10/09 10:14 miham 3rd HV change for BB Scintil. planes
  • 272311 05/10/09 09:54 R. Lindgren Happex start
  • 272308 05/10/09 09:20 miham dE/E plots after 2nd iteration of gain maching
  • 272307 05/10/09 08:50 R. Lindgren Beam back
  • 272305 05/10/09 08:49 R. Lindgren beam off
  • 272301 05/10/09 08:42 R. Lindgren Day shift starting
  • 272295 05/10/09 08:00 pking Owl shift summary.
  • 272294 05/10/09 07:56 Target target: Spin Flip Report: TO Report - OWL
  • 272285 05/10/09 06:55 pking Trouble starting Comptons runs 16877 & 16878; started Compton run 16879 at 06:50
  • 272280 05/10/09 05:56 miham Next HV changes for BB Scintil. planes (this time preformatted)
  • 272279 05/10/09 05:55 miham Next HV changes for BB Scintil. planes
  • 272276 05/10/09 05:45 miham BB mwdc are OFF
  • 272273 05/10/09 05:33 miham dE/E plots after 1st iteration of gain matching
  • 272272 05/10/09 05:27 miham Again Dead TDC channels in BB dE-plane
  • 272271 05/10/09 04:55 pking Started HAPPEX run 31355 at 04:50
  • 272270 05/10/09 04:54 pking Started Compton run 16876 at 04:47
  • 272265 05/10/09 04:41 miham Very slow BB analysis
  • 272264 05/10/09 04:18 pking spot++ raster check
  • 272250 05/10/09 03:32 miham BB dE/E plots before HV changes
  • 272245 05/10/09 03:09 miham Change of HV in some of the BB channels
  • 272240 05/10/09 02:22 pking Started Compton run 16875 at 02:11
  • Saturday

  • 272226 05/09/09 23:57 Xiaohui shift summary
  • 272225 05/09/09 23:56 Yi Qiang Experiment Summary for May 9, 2009
  • 272220 05/09/09 23:43 Jin Huang Target Spin State Conventions & Double Checks
  • 272219 05/09/09 23:33 Yi Qiang Experiment Summary for May 5 - 8, 2009
  • 272215 05/09/09 22:46 St. John ND Veto Left TDC's
  • 272213 05/09/09 22:36 adaq Hall A Tool
  • 272208 05/09/09 21:46 St. John ND N1 left TDC's run 20963
  • 272207 05/09/09 21:18 zhangyi He3 cell Moss is ready to go
  • 272202 05/09/09 20:21 hyde compton
  • 272188 05/09/09 19:18 miham Spot++ check before production
  • 272185 05/09/09 19:11 miham Spot++ check (#2053)
  • 272180 05/09/09 19:09 miham Spot++ check (#2052)
  • 272175 05/09/09 19:07 Yi Qiang Harp Scan result
  • 272172 05/09/09 19:04 Jin Huang Target -> He3
  • 272171 05/09/09 19:03 xiaohui raster check
  • 272166 05/09/09 18:57 Jin Huang Target -> BeO
  • 272165 05/09/09 18:50 miham BB HV turned OFF
  • 272164 05/09/09 18:49 bogdanw HRS-R VDC and HV supply
  • 272163 05/09/09 18:37 Jin Huang Target -> Empty
  • 272162 05/09/09 18:20 Target target: Spin Flip Report: Code Update; Restart Spin Flip Control VI
  • 272149 05/09/09 17:34 R. Lindgren Day Shift Summary
  • 272138 05/09/09 16:39 Jin Huang iocha5 reboot
  • 272137 05/09/09 16:31 R. Lindgren raster disaster
  • 272136 05/09/09 16:00 R. Lindgren Hall A access
  • 272135 05/09/09 15:52 miham Run #2043 is not a production run
  • 272134 05/09/09 15:46 miham BB is being turned ON
  • 272132 05/09/09 15:40 R. Lindgren VDC bottom high voltage alarm
  • 272130 05/09/09 15:35 miham One ND Channel was OFF
  • 272128 05/09/09 15:31 R. Lindgren VDC high voltage problem
  • 272126 05/09/09 15:14 M.Meziane R-arm VDC problems screenshots (part2)
  • 272125 05/09/09 15:12 M. Meziane -arm VDC problems screenshots (part 1)
  • 272119 05/09/09 14:07 R. Lindgren VDC problem
  • 272118 05/09/09 13:53 R. Lindgren VDC high voltage
  • 272117 05/09/09 13:48 R. Lindgren VDC high voltage down
  • 272116 05/09/09 13:31 A. Deur Compton DAQ using now prod_fa
  • 272112 05/09/09 12:43 M. Meziane check list: Hall A Tools
  • 272098 05/09/09 11:59 Yi Qiang MCC Raster Screw Up
  • 272097 05/09/09 11:58 R. Lindgren Production running
  • 272094 05/09/09 11:44 R. Lindgren Summary of events from 10:00 to 11:30
  • 272093 05/09/09 11:42 miham ND HV trip
  • 272083 05/09/09 11:28 Yi Qiang Raster size check for 4x4 after the raster-accident
  • 272080 05/09/09 11:15 A. Deur Ref cell vented
  • 272079 05/09/09 11:10 A. Deur Cell temperatures with beam hitting the wall.
  • 272076 05/09/09 10:58 Yi Qiang Raster check: 6x6
  • 272075 05/09/09 10:56 Yi Qiang Rates of 2uA beam on 3He target with 6x6 raste
  • 272072 05/09/09 10:53 Yi Qiang Raster size check for 4x4
  • 272070 05/09/09 10:52 Yi Qiang Rates of 2uA beam on 3He target with 4x4 raster
  • 272064 05/09/09 10:46 Moffit Re: DAQ problems
  • 272056 05/09/09 10:26 Kalyan L-HRS DAQ problems
  • 272045 05/09/09 10:06 miham Spot++ check
  • 272042 05/09/09 10:03 Yi Qiang Trigger Rates with 5.5uA and 4x4 raster
  • 272041 05/09/09 09:58 Kalyan 100% deadtime problem solved -- back to production
  • 272040 05/09/09 09:51 R. Lindgren Beam Position y
  • 272039 05/09/09 09:50 R. LIidgren DAQ Restored
  • 272038 05/09/09 09:45 Kalyan Runs 2017-2025 and 20932-20937 are junk
  • 272029 05/09/09 09:29 R.Lindgren DAQ Crashed
  • 272026 05/09/09 09:12 miham BB Scint. planes are ON
  • 272008 05/09/09 08:17 R. Lindgren Day shift starting
  • 272007 05/09/09 08:11 A. Deur Laser Absorption Spectrum
  • 272005 05/09/09 08:08 Target target: Spin Flip Report: spin flip system control report
  • 272004 05/09/09 08:05 R. Michaels request to use Prod_FA
  • 272003 05/09/09 07:58 pking Owl shift summary
  • 272002 05/09/09 07:49 ellie Set HAND Golden Run to 20928
  • 272001 05/09/09 07:44 Target target: Spin Flip Report: TO Report - OWL
  • 272000 05/09/09 07:35 ellie LHRS and RHRS CODAs Unhappy
  • 271988 05/09/09 06:41 ellie ROC10 Stopped responding, restarted LHRS CODA, rebooted ROC10, back working
  • 271984 05/09/09 06:32 Higinbotham Polarization After Pass Change
  • 271983 05/09/09 06:28 ellie Updated HAND Pedestals in DB
  • 271982 05/09/09 05:57 pking Started new HAPPEX run 31351 at 05:55
  • 271969 05/09/09 04:21 ellie Changed Golden Runs to 20924 (RHRS) and 2006 (LHRS)
  • 271968 05/09/09 04:16 ellie VDC Efficiency Screen Grabs
  • 271957 05/09/09 03:48 ellie LHRS CODA Crashed on Run 2005
  • 271948 05/09/09 02:58 Bradshaw Turn Auto-Flip back on
  • 271944 05/09/09 02:54 adaq Hall A tools
  • 271939 05/09/09 02:05 pking spot++ raster is 4x4
  • 271934 05/09/09 01:52 pking Increase event limit to 5M events on LHRS DAQ
  • 271929 05/09/09 01:37 pking Started new HAPPEX run 31350 at 01:34
  • 271916 05/09/09 00:50 pking He3 target is in request 2uA to check rates.
  • 271915 05/09/09 00:50 A. Camsonne Compton vertical scan
  • 271914 05/09/09 00:48 pking Raster check
  • 271913 05/09/09 00:47 Bradshaw Moved from Optics to 3He Target
  • 271910 05/09/09 00:33 Bradshaw Move from Empty to Optics
  • 271909 05/09/09 00:29 Bradshaw Moved from Reference to Empty
  • 271908 05/09/09 00:11 Bo/Yi Another check on RHRS dead time
  • 271907 05/09/09 00:04 Yi Qiang Reference cell is filled with H2 135 psig
  • 271905 05/09/09 00:00 Xiaohui Zhan shift summary
  • Friday

  • 271902 05/08/09 23:49 dwang Beam Polarization
  • 271879 05/08/09 22:26 Yi Qiang Reference cell raster size check: 5.5x6
  • 271874 05/08/09 22:18 Yi Qiang Reference cell raster size check: 4x4
  • 271871 05/08/09 22:15 Yi Qiang Raster Check for reference cell: 2.5x2
  • 271864 05/08/09 21:59 Yawei RHRS S1&S2 TDC plots
  • 271855 05/08/09 21:25 Jin Huang 20psig N2-> Ref cell
  • 271846 05/08/09 21:13 Jin Huang Target -> Ref Cell; Gauge Reading
  • 271841 05/08/09 20:24 xiaohui Beam is back!
  • 271840 05/08/09 20:16 xiaohui Beam OFF--MCC called and problems with RF
  • 271805 05/08/09 18:10 deng Lumi HV adjustment
  • 271800 05/08/09 17:40 St. John ND TDC's
  • 271799 05/08/09 17:37 Sirish Beam Polarization
  • 271764 05/08/09 16:27 Xiaohui Target moved to 3He
  • 271761 05/08/09 16:25 xiaohui raster size
  • 271758 05/08/09 16:20 xiaohui raster size check
  • 271755 05/08/09 16:18 Higinbotham BCM Calibration Check After Change
  • 271752 05/08/09 16:12 jpchen raster size 4x2.5(MCC)
  • 271749 05/08/09 16:08 jpchen Day shift summary
  • 271745 05/08/09 15:57 jpchen raster check
  • 271743 05/08/09 15:51 jpchen restarted coda for both HRS'
  • 271738 05/08/09 15:23 jpchen harp scan results fine
  • 271737 05/08/09 15:22 jpchen moved to BeO target
  • 271736 05/08/09 15:06 jpchen beam tuning, position and harp
  • 271735 05/08/09 14:40 jpchen ND HVs off for beam turning
  • 271734 05/08/09 14:37 jpchen HRS momentum change (both)
  • 271733 05/08/09 14:36 jpchen Hall back to beam permit and beam tuning now
  • 271732 05/08/09 14:12 Bo/Vince ROC8 busy singal check
  • 271731 05/08/09 14:02 R. Subedi/X. Access activity [PVDIS work]
  • 271730 05/08/09 14:00 jpchen target moved to EMPTY
  • 271729 05/08/09 13:53 Higinbotham BCM Filter 5V Supply Changed
  • 271726 05/08/09 12:48 jpchen pass change started at noon. Hall controlled access activities.
  • 271723 05/08/09 12:11 A. Camsonne Electron detector in garage position
  • 271707 05/08/09 11:12 jpchen Left Coda (ROC5) stuck, restarted, all fine.
  • 271705 05/08/09 11:07 Yi Qiang BCM Test with Compton Detectors OFF
  • 271703 05/08/09 10:58 A. Camsonne Electron detector back in data taking position
  • 271702 05/08/09 10:55 A. Camsonne Test Edet off and in garage
  • 271701 05/08/09 10:48 A. Camsonne Electron detector moved to garage position for test
  • 271700 05/08/09 10:45 Kalyan New HAPPEX run 31340 started
  • 271697 05/08/09 10:26 jpchen data taking status: smooth running
  • 271696 05/08/09 10:22 jpchen BCM screen shot: jumping problem still there
  • 271691 05/08/09 09:06 Kalyan Right arm VDC screen grab
  • 271686 05/08/09 08:51 Kalyan Left arm VDC screen grabs
  • 271685 05/08/09 08:25 A. Camsonne Compton configuration to Prod_FA starting run 16846
  • 271680 05/08/09 08:08 Target target: Spin Flip Report: TO Report - Owl
  • 271678 05/08/09 08:00 M Jones Owl shift summary
  • 271677 05/08/09 07:55 M Jones BCM going bad again
  • 271676 05/08/09 07:34 R. Michaels Compton FADC data
  • 271671 05/08/09 07:14 A. Camsonne Electron detector position from expert screen
  • 271670 05/08/09 07:00 ellie BCM is getting funky again
  • 271669 05/08/09 06:51 Alex Edet results, part III
  • 271664 05/08/09 06:46 Alex Edet results, part II
  • 271663 05/08/09 06:43 Alex Compton Edet results, part I
  • 271662 05/08/09 06:10 M Jones Hall A General Tools
  • 271661 05/08/09 06:07 M Jones SPOT ++ run 1950
  • 271656 05/08/09 05:29 A. Camsonne Compton Electron detector current
  • 271655 05/08/09 05:26 A. Camsonne Compton electron detector timing smaller time scale
  • 271637 05/08/09 03:58 ellie Efficiency Screen Grabs
  • 271636 05/08/09 03:56 ellie Started New HAPPEX Run (31338)
  • 271634 05/08/09 03:47 M Jones ROC5 crashed to end run 1946, Reboot ROC5
  • 271623 05/08/09 01:20 ellie RHRS Replay Screen "Total Shower E/P" Not Showing
  • 271618 05/08/09 00:22 A. Camsonne Compton run 16841 in Prod_FT
  • Thursday

  • 271612 05/07/09 23:57 Xiaohui Shift Summary
  • 271611 05/07/09 23:41 A. Camsonne Electron photon coincidence setup
  • 271610 05/07/09 23:21 V. Sulkosky Check for Run 20863
  • 271605 05/07/09 23:17 Yi Qiang LHRS is Buffered and RHRS is UnBuffered
  • 271604 05/07/09 23:09 Yi Qiang Plan for tonight's production runs and the pass change
  • 271603 05/07/09 22:54 Target target: Spin Flip Report: 05_07_09_swing_report
  • 271602 05/07/09 22:52 Yi Qiang Typo in the title, should be "Change of the BCM Downconverter"
  • 271601 05/07/09 22:50 Yi Qiang BCM Problems
  • 271600 05/07/09 22:42 V. Sulkosky/ Right HRS DAQ Checks
  • 271599 05/07/09 22:41 A. Camsonn Compton electron detector in beam position ( late entry )
  • 271590 05/07/09 22:25 A. Camsonne Electron detector in beam check
  • 271587 05/07/09 22:02 V. Sulkosky Re: multiple pulsing T1,T2
  • 271586 05/07/09 22:02 V. Sulkosky Re: multiple pulsing T1,T2
  • 271583 05/07/09 21:58 mcanan laser absorption spectrum
  • 271582 05/07/09 21:54 mcanan Target Motion
  • 271581 05/07/09 21:53 Hari spot++
  • 271578 05/07/09 21:38 mcanan Target Motion
  • 271573 05/07/09 20:36 R. Michaels multiple pulsing T1,T2
  • 271562 05/07/09 19:24 A. Camsonne Electron detector test plan
  • 271544 05/07/09 17:50 ycwang RUN 1935 and RUN 20847 are test runs.
  • 271542 05/07/09 17:42 Higinbotham Switch To Backup Down Converter
  • 271525 05/07/09 16:17 Higinbotham Shift Summary
  • 271524 05/07/09 16:15 A. Camsonne Current Compton Hall A Electron detector position calibration
  • 271523 05/07/09 16:03 miham ND HV trip
  • 271522 05/07/09 15:58 L. Zhu target: move to BeO and empty and back to He3
  • 271517 05/07/09 15:51 Higinbotham Raster Check
  • 271514 05/07/09 15:17 Bo Scaler check between L and R arms
  • 271513 05/07/09 15:10 V. Sulkosky/ R-HRS Deadtime Check
  • 271512 05/07/09 14:39 Yawei Grading for the Fiber Tips
  • 271509 05/07/09 14:36 Yawei Grading for the Fiber Tips
  • 271505 05/07/09 14:12 Higinbotham BCM Problem Solved?
  • 271495 05/07/09 13:44 V. Sulkosky R-HRS Raster X Underflow
  • 271494 05/07/09 13:38 St. John ND N1 Rgt TDC's
  • 271493 05/07/09 13:34 Higinbotham Techs Are Checking for Spares
  • 271488 05/07/09 13:17 V. Sulkosky Added Shower Display to Online Replay
  • 271487 05/07/09 13:14 Higinbotham Found Spare 10 KHz BCM
  • 271486 05/07/09 13:13 V. Sulkosky R-HRS Preshower and Shower Pedestals
  • 271485 05/07/09 13:10 St. John Corrections for start of run entry
  • 271484 05/07/09 13:06 Moffit Re: HAPPEX BCM
  • 271479 05/07/09 13:00 Moffit HAPPEX view of BCMs.. is the problem with the downconverter?
  • 271474 05/07/09 12:29 St. John Runs 1925 & 20833
  • 271473 05/07/09 12:27 Higinbotham Right Arm Shower & PreShower Ped. Updated
  • 271451 05/07/09 11:16 Kalyan Added NormAna to R-HRS replay
  • 271450 05/07/09 11:05 A. Camsonne Asymmetry electron detector run 16758
  • 271449 05/07/09 11:01 Higinbotham Shower & PreShower Pedistals
  • 271448 05/07/09 10:59 Higinbotham CODA Right Only Find High Deadtime Still
  • 271447 05/07/09 10:57 St. John ND N1 Lft TDC's run 20830
  • 271442 05/07/09 10:28 Target target: Spin Flip Report: No Shift Worker Comment
  • 271441 05/07/09 10:27 Higinbotham Hall A BCM Change Seen
  • 271440 05/07/09 10:25 L. Zhu laser absorption spectrum
  • 271439 05/07/09 10:16 St. John New happex run 31336
  • 271436 05/07/09 10:02 St. John ND N1 Right TDC's
  • 271435 05/07/09 09:43 L. Zhu target: move to BeO and back to He3
  • 271434 05/07/09 09:36 Moffit/Mihov Rebooted ROC26
  • 271433 05/07/09 09:15 Yi Qiang Hall in Controlled Access for Rebooting ROC26
  • 271428 05/07/09 08:41 Kalyan R-HRS online VDC plots
  • 271424 05/07/09 08:03 V. Sulkosky Re: Still have problems with BCMs
  • 271423 05/07/09 08:03 M Jones Owl shift summary
  • 271422 05/07/09 07:58 V. Sulkosky Re: L-arm DAQ run start/end windows doesn't pop up
  • 271421 05/07/09 07:46 hoyoung L-arm DAQ run start/end windows doesn't pop up
  • 271416 05/07/09 07:24 M Jones Attached GIF figure of BCM
  • 271415 05/07/09 07:20 M Jones Still have problems with BCMs
  • 271414 05/07/09 07:04 M Jones After restarting left arm CODA, Deadtime is now 13% for left
  • 271413 05/07/09 07:02 M Jones Left arm CODA hangs for runs 1917-1918
  • 271398 05/07/09 05:25 M Jones Hall A tools screen
  • 271397 05/07/09 05:22 M Jones spot++ run 1915
  • 271392 05/07/09 04:26 ellie Right Arm Efficiencies Screen Grab
  • 271391 05/07/09 04:22 ellie Left Arm Efficiency Screen Capture
  • 271386 05/07/09 04:13 M Jones ROC 26 is hung up, cannot take Lumi data
  • 271362 05/07/09 01:20 ellie Changed LHRS Golden Run to 1908
  • 271359 05/07/09 01:15 ellie Typo in Run Number Re: Problem with hac_bcm_average
  • 271358 05/07/09 01:12 M Jones Started new HAPPEX run 31334 at 00:30
  • 271357 05/07/09 01:12 ellie Changed RHRS/HAND Golden Run to 20816
  • 271356 05/07/09 00:48 M Jones BCM scalers also so a shift
  • 271355 05/07/09 00:45 hoyoung ND plots changed
  • 271350 05/07/09 00:22 M Jones Problem with hac_bcm_average
  • 271349 05/07/09 00:06 A. Puckett Swing shift summary
  • Wednesday

  • 271343 05/06/09 23:37 A. Deur laser absorption spectrum.
  • 271342 05/06/09 23:36 Target target: Spin Flip Report: Spin flip system report
  • 271341 05/06/09 23:30 V. Sulkosky Re: R-HRS Deadtime with buffered Mode
  • 271340 05/06/09 23:21 Yi Qiang Trigger Rate with 10uA beam
  • 271339 05/06/09 23:20 golge Run #1907 L-arm VDC
  • 271338 05/06/09 23:15 zhangyi RHRS dead time from run 20814
  • 271337 05/06/09 23:14 Yi Qiang R-HRS is running with Unbuffered Mode and Neutron Detector ON
  • 271332 05/06/09 22:35 golge shift in plot R-arm S1 ADC2
  • 271329 05/06/09 22:20 puckett spot++ looks good, back to 10 uA
  • 271326 05/06/09 22:19 golge spot++ check Run#1905
  • 271325 05/06/09 22:18 puckett MCC sends CW-->asked for 2uA for spot++
  • 271322 05/06/09 22:13 puckett Started production--then orbit problems-->pulsed beam
  • 271317 05/06/09 21:55 puckett Harp scan successful
  • 271312 05/06/09 21:43 A. Deur Target polarization 65.7%
  • 271309 05/06/09 21:40 golge spot++ check Run#1902 6x6
  • 271308 05/06/09 21:39 golge scaler check #1902
  • 271303 05/06/09 21:33 golge t3 rate check run#1901
  • 271302 05/06/09 21:28 puckett BPMs with CW on 3He
  • 271297 05/06/09 21:21 golge spot++ Run#1900
  • 271294 05/06/09 21:17 puckett Hall B informs us that they will change the IHWP status to OUT!
  • 271293 05/06/09 21:05 puckett Beam is back, 2 uA on BeO looks good, moved to empty target for harp scan
  • 271292 05/06/09 20:57 V. Sulkosky re: roc8 busy
  • 271289 05/06/09 20:14 brads Happex GUI restarted on vncviewer.
  • 271288 05/06/09 20:12 R. Michaels re: roc8 busy
  • 271278 05/06/09 19:57 brads killed/restarted X server on adaql6
  • 271273 05/06/09 18:42 R. Michaels re: roc8 busy
  • 271270 05/06/09 16:13 A. Deur Target is in C12 position
  • 271269 05/06/09 16:12 Techs Neutron Detector status
  • 271266 05/06/09 16:07 spiegel Argon/ethane 50/50 mix
  • 271263 05/06/09 16:00 R. Subedi Shift Summary
  • 271258 05/06/09 15:33 V. Sulkosky Re: L-HRS Gas Cerenkov ADC # 6
  • 271229 05/06/09 13:51 Yawei Swing shift summary(May 5th)
  • 271228 05/06/09 13:51 Moffit ROC8 Busy to TS now functional.
  • 271227 05/06/09 13:47 Yawei Swing shift summary(May 5th)
  • 271218 05/06/09 13:19 Moffit/Subed ROC8 Busy to RightHrs TS wasn't working.
  • 271217 05/06/09 13:14 R. Michaels R-HRS: TS program change
  • 271205 05/06/09 11:57 katich Lasers back on
  • 271204 05/06/09 11:54 Yi Qiang Target Survey Done
  • 271203 05/06/09 10:24 gjin gas shed: argon/ethane one bottle empty
  • 271202 05/06/09 10:09 R. Subedi Hall in controlled access since 08:30
  • 271201 05/06/09 08:53 R. Michaels DAQ change last night (LATE entry)
  • 271200 05/06/09 07:42 Target target: Spin Flip Report: TO Report - Owl
  • 271198 05/06/09 07:34 M Jones owl shift summary
  • 271197 05/06/09 06:06 M Jones Hall A Tools page, Right ARM dipole finally set
  • 271187 05/06/09 01:15 M Jones Hall access finished
  • 271186 05/06/09 01:04 M Jones Left HRS is up
  • 271185 05/06/09 00:55 M Jones Hall access
  • 271184 05/06/09 00:34 Bradshaw Return of the Lasers
  • 271183 05/06/09 00:21 Bradshaw Laser Trip
  • 271182 05/06/09 00:11 M Jones tools page
  • 271180 05/06/09 00:02 golge hall a tools
  • Tuesday

  • 271179 05/05/09 23:45 Yawei Hall A tools
  • 271178 05/05/09 23:38 golge cryo levels
  • 271177 05/05/09 23:37 golge hall a tools
  • 271176 05/05/09 23:29 chench target condition after the CHL down
  • 271175 05/05/09 23:28 chench target status after the CHL down
  • 271174 05/05/09 23:23 golge cryogen levels
  • 271173 05/05/09 23:22 golge hall a tools snapshot
  • 271172 05/05/09 23:16 Yawei End the Lumi run
  • 271171 05/05/09 23:07 golge Beam Down at 23:02
  • 271170 05/05/09 23:07 chench turn off the auto flip
  • 271167 05/05/09 22:57 golge spot++ Run # 1886 4x4
  • 271164 05/05/09 22:52 chench Move the target to from reference cell to 3He
  • 271161 05/05/09 22:28 adaq test
  • 271159 05/05/09 22:27 R.Michaels halog test
  • 271158 05/05/09 22:20 R. Michaels Compton test runs
  • 271155 05/05/09 21:35 yawei reference cell: H2, 135 psig trigger rates
  • 271154 05/05/09 21:27 chench Move target from 3He to reference cell, with pressure of H2 = 135psig
  • 271149 05/05/09 21:20 golge spot++ 4x4 Run #1883
  • 271144 05/05/09 21:05 Yawei Lumi issues
  • 271143 05/05/09 20:49 golge R-arm plot shift #20789
  • 271142 05/05/09 20:30 golge Neutron Plot Run # 20788
  • 271137 05/05/09 20:06 golge Lumi
  • 271119 05/05/09 19:42 Bo Re:Access activity for T8 problem
  • 271116 05/05/09 19:38 chench Move the target from optics to 3He
  • 271113 05/05/09 19:33 Yawei scaler BeO, 5uA run(#1878) and(#20785)
  • 271108 05/05/09 19:29 golge spot++ for Run # 1877
  • 271101 05/05/09 19:13 R. Subedi Access activity for T8 problem
  • 271100 05/05/09 19:10 chench move target from 3He to optics
  • 271079 05/05/09 18:08 Yawei Controlled Access
  • 271070 05/05/09 17:50 R. Michaels re: deadtime story
  • 271043 05/05/09 17:06 Moffit RightHRS deadtime ~45%, HAND Busy suspected..
  • 271033 05/05/09 16:39 Yawei T3 check
  • 271031 05/05/09 16:37 chench move the target from optics to 3He, and flit the spin
  • 271024 05/05/09 16:22 Yawei spot++, 6x6 BeO
  • 271019 05/05/09 16:12 Yawei spot++ 2x2
  • 271014 05/05/09 16:04 Yawei spot++ ,4x4
  • 271010 05/05/09 16:00 Holmstrom Day Shfit Summary
  • 271009 05/05/09 15:35 Techs/spiege Detector Hut air conditioners
  • 271008 05/05/09 15:28 Sirish Compton Electron Detector
  • 271007 05/05/09 15:26 Bo cratemap for scaler replay of R-arm runs
  • 271006 05/05/09 15:09 gjin target moved to carbon
  • 271005 05/05/09 14:18 Ole Hansen C++ Analyzer updated
  • 271004 05/05/09 13:51 Ramesh/Xiaoc Access Activity (changes in PVDIS setup only)
  • 271003 05/05/09 13:49 miham BB dE dead channel #9 fixed
  • 271002 05/05/09 13:45 Ewing/Fansle Lead wall
  • 271001 05/05/09 13:27 Yi Qiang NMR Cross Calibration
  • 270999 05/05/09 13:11 Yi Qiang Internal Temperature Test
  • 270998 05/05/09 12:30 Moffit/Mihov BB Triggers now into LeftHRS scalers.
  • 270997 05/05/09 12:24 Moffit TECHS: LeftARM AC unit dripping water.
  • 270996 05/05/09 12:19 Yi Qiang EPR Calibration for Vertical Pumping
  • 270986 05/05/09 11:59 Yawei Cuts after modifying
  • 270985 05/05/09 11:43 R. Michaels Compton test runs.
  • 270982 05/05/09 11:25 R. Michaels Compton unified DAQ
  • 270981 05/05/09 11:24 Bo/Yawei Cable connection check for ADC #6 of L-arm Gas Cerenkov
  • 270980 05/05/09 11:03 Yi Qiang Gauges for reference cell have been zeroed
  • 270976 05/05/09 10:38 Yi Qiang BigBite is ON now
  • 270974 05/05/09 09:53 Yi Qiang Field ajusted for BigBite OFF and target moved to pickup coils for internal temperature test
  • 270973 05/05/09 09:51 Yi Qiang BigBite is OFF
  • 270972 05/05/09 08:17 gjin target auto flip suspended
  • 270971 05/05/09 08:14 R. Subedi Re: R-arm Preshower ADC 18 problem
  • 270970 05/05/09 07:36 Target target: Spin Flip Report: Target Operator Report: Owl
  • 270969 05/05/09 07:20 miham LUMI replay plots for the OWL shift runs
  • 270968 05/05/09 07:06 M Jones Owl shift summary
  • 270966 05/05/09 07:02 M Jones Radcon in process of hall survey
  • 270964 05/05/09 07:01 M Jones Beam off for beam studies
  • 270962 05/05/09 06:40 M Jones spot++ for run 1853
  • 270957 05/05/09 06:14 miham Modified LUMI Replay scripts
  • 270956 05/05/09 05:56 M Jones Hall A tools screen
  • 05/05/09 05:24 miham Suspicious LUMI Plots for run #1845
  • 270943 05/05/09 04:08 miham New Compton run #16803
  • 270933 05/05/09 03:51 miham ND HV trip
  • 270924 05/05/09 02:17 M Jones Start HAPPEX run 31328
  • 270919 05/05/09 02:04 miham New Compton Run #16802
  • 270914 05/05/09 00:41 hoyoung R-arm Preshower ADC 18 problem
  • 270909 05/05/09 00:04 Ole Hansen Swing Shift Summary
  • Monday

  • 270907 05/04/09 23:46 golge Spot ++ Run# 1844
  • 270902 05/04/09 22:54 golge Junk Data Due to Beam trips
  • 270899 05/04/09 22:45 Ole Hansen Beam back. HV trips
  • 270898 05/04/09 22:13 Zhihong Ye Target Graph 2
  • 270897 05/04/09 22:10 Zhihong Ye screen shot for target spactra
  • 270896 05/04/09 22:04 Ole Hansen New HAPPEX run
  • 270895 05/04/09 20:51 Ole Hansen adaql2 X freeze
  • 270894 05/04/09 20:49 Ole Hansen Beam down
  • 270872 05/04/09 19:08 Yawei The cuts I used for calculation
  • 270871 05/04/09 19:07 Yawei The cuts I used for calculation
  • 270851 05/04/09 16:24 Ole Hansen New HAPPEX run
  • 270848 05/04/09 16:16 averett day shift summary
  • 270842 05/04/09 15:30 Moffit Re: Compton... it's fine now.
  • 270841 05/04/09 15:29 Moffit Re: HAND HV change.
  • 270840 05/04/09 15:22 adaq lumi check: normal
  • 270839 05/04/09 15:02 gjin compton: is it wrong?
  • 270838 05/04/09 14:58 Moffit HAND HV change for several channels... for 20740
  • 270827 05/04/09 13:16 chench Move target from Reference Cell to 3He cell
  • 270822 05/04/09 12:39 averett new HAPPEX run started
  • 270817 05/04/09 12:05 averett D2 in reference cell
  • 270812 05/04/09 11:55 chench Move target from optics to Reference Cell
  • 270786 05/04/09 10:44 averett Reference cell H2 changed to D2
  • 270785 05/04/09 10:37 chench Move target from 3He to optics
  • 270784 05/04/09 10:37 gjin start compton run 16794
  • 270783 05/04/09 10:36 gjin compton run 16793: end of run failed
  • 270780 05/04/09 09:46 Moffit Added All Neutron Detector thresholds to Left and Right Start of Run HALOG entries.
  • 270779 05/04/09 09:41 Moffit Re: BB triggers
  • 270774 05/04/09 09:06 chench Flip the spin of the t arget
  • 270773 05/04/09 09:02 gjin missing entry of T1, T2, and T5 rates
  • 270772 05/04/09 08:43 Moffit Re: BB triggers in End-of-Run
  • 270769 05/04/09 08:37 averett new HAPPEX run 31324
  • 270768 05/04/09 08:35 averett Beam half wave plate IN
  • 270767 05/04/09 08:31 averett new HAPPEX run
  • 270763 05/04/09 08:13 miham LUMI replay plots for OWL shift runs
  • 270762 05/04/09 08:02 Target target: Spin Flip Report: To Report: Owl
  • 270761 05/04/09 08:00 Bradshaw Shift Leader Report: Owl
  • 270760 05/04/09 07:59 Bradshaw Spot++ Run:1822
  • 270751 05/04/09 06:38 miham New Compton run#16793
  • 270750 05/04/09 06:23 Bradshaw Hall A Tools Screen
  • 270749 05/04/09 06:21 Bradshaw Spot++ Run: 1820
  • 270739 05/04/09 04:37 Bradshaw Happex Restart: 31322
  • 270735 05/04/09 04:34 miham New Compton run #16792
  • 270734 05/04/09 03:54 Bradshaw Spot++ Run: 1817
  • 270725 05/04/09 02:54 miham Question about Trigger rates in the End-of-run file?
  • 270724 05/04/09 02:37 miham New Compton run #16791
  • 270723 05/04/09 02:18 Bradshaw Spot ++ Run:1815
  • 270714 05/04/09 00:38 Bradshaw Spot++ Run:1813
  • 270711 05/04/09 00:06 Target target: Spin Flip Report: TO Report - Swing
  • 270710 05/04/09 00:04 LeRose Shift Summary (swing)
  • 270709 05/04/09 00:02 LeRose Shift Summary (swing)
  • 270708 05/04/09 00:01 Yawei Start a new happex run#31321
  • Sunday

  • 270706 05/03/09 23:36 LeRose No Beam
  • 270695 05/03/09 22:17 LeRose Hall A Tools
  • 270694 05/03/09 22:13 LeRose Spot++
  • 270685 05/03/09 20:11 Yawei New happex run#31320
  • 270683 05/03/09 20:09 Yawei New compton run#16787
  • 270675 05/03/09 18:41 V. Sulkosky Left HRS Gas Cerenkov #6
  • 270673 05/03/09 18:38 LeRose Beam Back
  • 270671 05/03/09 18:22 LeRose No Beam
  • 270669 05/03/09 18:19 Yawei Neutron Detector HV tripped.
  • 270663 05/03/09 17:49 Yawei Lumi replay results, Run#1803
  • 270662 05/03/09 17:33 Yawei New compton run started at 17:32, run#16787
  • 270661 05/03/09 17:30 Yawei Plots of Left arm Gas Cerenkov, ADC 6
  • 270660 05/03/09 17:25 Yawei Seems that so called "Self Healing" only worked less than 20 runs
  • 270655 05/03/09 16:37 R. Michaels MSS problems fixed
  • 270654 05/03/09 16:37 gjin day shift summary
  • 270640 05/03/09 15:06 R. Michaels problems writing data to MSS
  • 270639 05/03/09 14:51 R. Michaels cannot revive FADC crate
  • 270638 05/03/09 14:50 R. Michaels it was a mistake
  • 270637 05/03/09 14:46 gjin DAQ complaining about disk running out of space
  • 270626 05/03/09 14:04 gjin lumi check
  • 270625 05/03/09 13:15 Zhihong Ye screen shot for target spactra
  • 270612 05/03/09 11:26 gjin beam tripped between 8:22 am and 9:11 am
  • 270611 05/03/09 11:05 gjin lumi check
  • 270610 05/03/09 10:59 gjin compton run 16783: end of run failed
  • 270609 05/03/09 10:54 R. Michaels request for a ROC reboot
  • 270604 05/03/09 10:37 gjin gas shed: carbon dioxide one bottle empty
  • 270599 05/03/09 09:40 Hari tools screen
  • 270598 05/03/09 09:25 gjin spot++ after a beam trip of 50 mins
  • 270595 05/03/09 08:42 miham Lumi plots for all runs taken during OWL shift
  • 270592 05/03/09 08:15 miham OWL Shift Summary (05/03/09)
  • 270590 05/03/09 07:55 miham Spot++ check (#1791)
  • 270581 05/03/09 06:54 miham New Compton run #16782
  • 270580 05/03/09 06:48 miham Self Healing Gas Cerenkov
  • 270579 05/03/09 06:18 miham Scaler snap shots
  • 270578 05/03/09 06:17 miham Spot++
  • 270569 05/03/09 04:54 adaq There is someproblem about the replay process
  • 270564 05/03/09 04:30 miham New Compton run #16781
  • 270563 05/03/09 04:04 miham New Happex run #31316 and LUMI plots
  • 270562 05/03/09 03:58 adaq Deadtime for both R-HRS and L-HRS
  • 270557 05/03/09 03:53 miham Scaler Snap Shots
  • 270556 05/03/09 03:48 miham Spot++ check (#1785)
  • 270551 05/03/09 03:06 miham More than one hour of bad beam
  • 270550 05/03/09 02:27 miham New Compton run #16780
  • 270549 05/03/09 02:15 miham Bad Beam
  • 270544 05/03/09 01:25 miham Scaler snap shots
  • 270543 05/03/09 01:23 miham Spot++ check (#1783)
  • 270542 05/03/09 01:20 miham New Happex run #31315 and LUMI plots
  • 270537 05/03/09 01:06 miham Temperatures of the fiber connections in the Laser Optics Enclosure
  • 270536 05/03/09 00:34 miham Compton problems and new Comton run #16779
  • 270531 05/03/09 00:17 zhangyi gain problem?
  • 1 05/03/09 00:07 averett Swing Shift Summary (05/02/09)
  • Saturday

  • 270529 05/02/09 23:52 miham Lumi replay for run #1777
  • 270520 05/02/09 22:40 Yawei Left arm Cerenkov ADC 6 looks not so good
  • 270515 05/02/09 22:03 miham New Compton run #16777
  • 270514 05/02/09 22:01 Yawei Good electrons we have since we changed to 3 pass
  • 270511 05/02/09 21:05 averett beam halfwave plate OUT
  • 270504 05/02/09 20:04 K Wang Hall A tool
  • 270503 05/02/09 19:35 averett 3He polarization decrease to 45.7%
  • 270500 05/02/09 19:30 averett pol3he target in beam
  • 270499 05/02/09 19:28 averett reference cell now empty
  • 270492 05/02/09 18:25 Yawei fill H2 in reference cell
  • 270481 05/02/09 17:36 averett N2 in reference cell
  • 270472 05/02/09 16:41 Yawei spot++ check, 2uA
  • 270467 05/02/09 16:31 Yawei Pressure 3He gas: 128 psig
  • 270466 05/02/09 16:29 Yawei Move target to refrence cell.
  • 270463 05/02/09 16:15 brads Charge Asym is OK now
  • 270461 05/02/09 16:01 A. Deur Day Shift summary
  • 270460 05/02/09 15:45 A. Deur Hall A tools
  • 270455 05/02/09 15:18 brads Re: Charge asymmetry is degrading.
  • 270454 05/02/09 15:11 brads Charge asymmetry is degrading.
  • 270453 05/02/09 15:01 A. Deur Happex DAQ
  • 270424 05/02/09 11:18 R. Michaels work on Compton DAQ
  • 270407 05/02/09 08:35 R. Michaels re: connection refused
  • 270406 05/02/09 08:31 miham LUMI Replay added to main replay Scripts
  • 270405 05/02/09 08:09 A. Deur laser absorption spectrum.
  • 270404 05/02/09 08:08 Target target: Spin Flip Report: Spin Flip system status
  • 270402 05/02/09 08:02 miham OWL Shift Summary (05/02/09)
  • 270397 05/02/09 07:33 miham Scaler Snap Shots
  • 270396 05/02/09 07:24 miham Spot++ check (#1757)
  • 270391 05/02/09 06:47 miham New Compton run #16770
  • 270390 05/02/09 06:44 miham DAQ problems
  • 270378 05/02/09 05:05 miham Scaler snap shots
  • 270377 05/02/09 05:02 miham Spot++ check (#1753)
  • 270374 05/02/09 04:46 miham New Happex run #31310 and LUMI PLOTS
  • 270371 05/02/09 04:36 miham HAPPEX Start/Stop Error during runs #1747 #20653
  • 270370 05/02/09 04:23 miham BB Dead TDC channel in dE plane
  • 270361 05/02/09 04:09 miham New Compon run #16769
  • 270356 05/02/09 03:24 miham A Change in the ND?
  • 270351 05/02/09 02:44 miham Scaler snap shots
  • 270350 05/02/09 02:43 miham Spot++ check (#1750)
  • 270345 05/02/09 02:13 miham LUMI Replays for runs #1748 and #20654
  • 270344 05/02/09 02:06 miham New Compton run #16768
  • 270339 05/02/09 01:16 miham Scaler snap shots
  • 270338 05/02/09 01:01 miham Spot++ check (#1748)
  • 270335 05/02/09 00:46 miham New Happex run #31309 and LUMI PLOTS
  • 270332 05/02/09 00:25 Bo Redundant scaler check on L-arm and R-arm
  • 270331 05/02/09 00:21 Moffit/Mihov Swing Shift Summary (05/01/09)
  • Friday

  • 270327 05/01/09 23:56 miham Compton Problem
  • 270324 05/01/09 23:52 miham Compton Problems
  • 270323 05/01/09 23:38 adaq Target
  • 270318 05/01/09 22:32 miham Replayed LUMI runs #1744 and #20650
  • 270289 05/01/09 18:55 Moffit Ugh, another BCM dx3 problem found
  • 270288 05/01/09 18:27 Moffit BCM dx3: removed patch cable... problem goes away.
  • 270285 05/01/09 18:18 Moffit Replaced a patch cable for BCM dx3 on CH Level translator
  • 270280 05/01/09 17:45 Moffit Beam returns... Resume production
  • 270279 05/01/09 17:17 Bo New HV settting on Preshower18 of RHRS
  • 270278 05/01/09 17:09 Moffit MCC: No beam due to PSS malfunction
  • 270271 05/01/09 16:07 R. Subedi Shift Summary
  • 270269 05/01/09 15:58 R. Subedi Beam back, taking production data
  • 270268 05/01/09 15:57 R. Subedi Runs with BB HV ON
  • 270265 05/01/09 15:41 R. Subedi BB HV off
  • 270264 05/01/09 15:37 R. Subedi Beam away
  • 270261 05/01/09 15:16 R. Subedi Raster scan rate result
  • 270260 05/01/09 15:04 R. Subedi Spot check on 3He
  • 270251 05/01/09 14:55 R. Subedi Spot check on 6x6
  • 270242 05/01/09 14:46 R. Subedi BB HV on while spot checking
  • 270241 05/01/09 14:44 R. Subedi Spot++ on 3He for 2x2
  • 270236 05/01/09 14:18 R. Subedi spot check for 2x2
  • 270231 05/01/09 14:09 R. Subedi spot++ for 6x6
  • 270226 05/01/09 14:00 R. Subedi Spot++ for 4x4
  • 270223 05/01/09 13:52 R. Subedi Harp result ok
  • 270222 05/01/09 13:33 gjin lumi check
  • 270221 05/01/09 13:27 R. Subedi Beam coming
  • 270220 05/01/09 13:26 R. Subedi access work (late entry)
  • 270217 05/01/09 13:21 Bo Presher18 in RHRS
  • 270208 05/01/09 11:49 R. Michaels access activity
  • 270207 05/01/09 11:49 R. Subedi Target changed to Optics target (Carbon)
  • 270206 05/01/09 11:44 Katich Lumi Replay, again
  • 270204 05/01/09 11:29 atrig Trigger-Download
  • 270201 05/01/09 11:24 Katich Lumi Data Replay
  • 270193 05/01/09 11:00 Moffit Returned BCM dx3 to normal
  • 270192 05/01/09 10:56 Moffit HAPPEX run for helicity v2f dx3
  • 270191 05/01/09 10:54 Moffit Helicity v2f p.2
  • 270189 05/01/09 10:50 Moffit Helicity v2f result p.1
  • 270186 05/01/09 10:44 Moffit Plugged in v2f helicity into dx3
  • 270185 05/01/09 10:43 R. Michaels R-HRS deadtime checks
  • 270161 05/01/09 09:58 Ramesh Asked for controlled access
  • 270149 05/01/09 08:07 miham New Compton run #16765
  • 270147 05/01/09 08:03 Bradshaw Shift Leader Summary - Owl
  • 270146 05/01/09 07:59 miham New Happex run #31302 started + LUMI PLOTS
  • 270144 05/01/09 07:48 Target target: Spin Flip Report: TO Report - Owl
  • 270140 05/01/09 07:01 Bradshaw Spot++ Check Run:1721
  • 270136 05/01/09 06:59 miham Results of the compton run #16763
  • 270130 05/01/09 06:07 miham New Compton run #16764
  • 270129 05/01/09 06:04 miham Compton Strip charts
  • 270124 05/01/09 04:46 miham Results of the compton run #16762
  • 270119 05/01/09 04:17 Sirish Compton Electron Spectrum
  • 270118 05/01/09 04:15 miham New Compton run #16763
  • 270117 05/01/09 04:13 miham New golden runs #1716 and #20610
  • 270108 05/01/09 04:06 Bradshaw Tools Screenshot
  • 270107 05/01/09 03:51 miham Started new Happex run #31301 + LUMI PLOTS
  • 270106 05/01/09 03:39 Bradshaw Rates For L/R-HRS
  • 270105 05/01/09 03:37 Bradshaw Trigger Rates
  • 270104 05/01/09 03:37 Bradshaw Spot++
  • 270099 05/01/09 03:13 miham Compton strip charts
  • 270094 05/01/09 02:24 Bradshaw Beam trips at 10 uA
  • 270093 05/01/09 02:10 miham New Compton run #16762
  • 270091 05/01/09 02:00 miham BB HV turned off again.
  • 270089 05/01/09 01:56 Sirish Compton Beam View
  • 270086 05/01/09 01:53 Bradshaw Spot++ - 6x6 Raster T3
  • 270085 05/01/09 01:47 Bradshaw Spot++ Check - 6x6 Raster
  • 270080 05/01/09 01:37 Bradshaw Spot++ Check- 2x2 Raster
  • 270079 05/01/09 01:32 W. Luo Reboot ts0 before 20607
  • 270076 05/01/09 01:15 Bradshaw Moved from Optics to He3
  • 270075 05/01/09 01:11 Bradshaw Moved from Optics to He3 and Back to Optics
  • 270071 05/01/09 00:20 miham BB HV on during calibration
  • 270070 05/01/09 00:17 Bradshaw Spot++ for Run 1712